### v2

# CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning

Abstract
Post-training compression of large language models (LLMs) largely relies on low-rank weight approximation, which represents each column of a weight matrix in a shared low-dimensional subspace. While this is a computationally efficient strategy, the imposed structural constraint is rigid and can lead to a noticeable model accuracy drop. In this work, we propose CoSpaDi (Compression via Sparse Dictionary Learning), a novel training-free compression framework that replaces low-rank decomposition with a more flexible structured sparse factorization in which each weight matrix is represented with a dense dictionary and a column-sparse coefficient matrix. This formulation enables a union-of-subspaces representation: different columns of the original weight matrix are approximated in distinct subspaces spanned by adaptively selected dictionary atoms, offering greater expressiveness than a single invariant basis. Crucially, CoSpaDi leverages a small calibration dataset to optimize the factorization such that the output activations of compressed projection layers closely match those of the original ones, thereby minimizing functional reconstruction error rather than mere weight approximation. This data-aware strategy preserves better model fidelity without any fine-tuning under reasonable compression ratios. Moreover, the resulting structured sparsity allows efficient sparse–dense matrix multiplication and is compatible with post-training quantization for further memory and latency gains. We evaluate CoSpaDi across multiple Llama and Qwen models under per-layer and per-group settings at 205020-50% compression ratios, demonstrating consistent superiority over state-of-the-art data-aware low-rank methods both in accuracy and perplexity. Our results establish structured sparse dictionary learning as a powerful alternative to conventional low-rank approaches for efficient LLM deployment.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of applications, from conversational agents Brown et al. (2020); OpenAI (2023) to general-purpose reasoning systems Touvron et al. (2023a); Anil et al. (2023), owing to their ability to capture long-range dependencies through attention mechanisms Vaswani et al. (2017); Devlin et al. (2019). However, this success entails substantial memory and computational demands during both training and inference, which severely limits their deployment in resource-constrained environments.
Various compression and acceleration techniques have been proposed over recent years to mitigate these challenges, including pruning Frankle & Carbin (2019); Sanh et al. (2020), quantization Dettmers et al. (2022); Xiao et al. (2023); Yao et al. (2022); Dettmers et al. (2023), and weight decomposition/factorization Denton et al. (2014); Hu et al. (2022); Liu et al. (2024); Zhang et al. (2023); Ma et al. (2019); Chen et al. (2018); Hsu et al. (2022); Yu & Wu (2023); Wang et al. (2025a); Han et al. (2015).
Recently, there has been growing interest in effective compression of large models without retraining or fine-tuning, since this is often computationally prohibitive. A prominent line of work leverages low-rank approximations of weight matrices via Singular Value Decomposition (SVD). Early approaches such as DRONE Chen et al. (2021), FWSVD Hsu et al. (2022) and ASVD Yuan et al. (2023) introduced activation-aware truncation of singular values, while SVD-LLM Wang et al. (2025a) and its variant SVD-LLMv2 Wang et al. (2025b) proposed truncation-aware and optimized singular value selection strategies. In parallel, Basis Sharing Wang et al. (2024) explored cross-layer sharing of a common basis to further enhance compression efficiency by exploiting inter-layer redundancy. Despite their effectiveness, these methods are inherently constrained by the use of a single shared low-dimensional subspace, which may limit representational flexibility and prevent full exploitation of redundancies in LLM parameters.
In this work, we argue that moving beyond the low-rank weight decomposition is a promising new direction. Specifically, we propose to adopt dictionary learning, a well-established paradigm in signal and image processing Aharon et al. (2006); Elad (2010), and apply it to LLM compression. Unlike low-rank approximations that confine all columns of a weight matrix to a shared linear subspace, dictionary learning enables a richer union-of-subspaces representation: each column is approximated using only a sparse subset of atoms from a learned dictionary, allowing different columns to reside in distinct subspaces spanned by different atom combinations. Such a representation better accommodates heterogeneous features and introduces additional flexibility to reduce the overall approximation error.
To this end, we introduce CoSpaDi (Compression via Sparse Dictionary Learning), a training-free framework that applies dictionary learning to jointly estimate dictionaries and sparse coefficients used for compressing LLM weight matrices. Our contributions are as follows:
  • We introduce dictionary learning with sparse coding as a new paradigm for LLM compression, addressing the limitation of using a single invariant basis for the weight approximation which is inherent in SVD-based methods.
  • We demonstrate that CoSpaDi is effectively integrated with data-aware optimization, yielding state-of-the-art compression performance for a wide range of compression ratios, while being compatible with post-training quantization for further memory and latency gains.
  • Through extensive experiments, we show that our approach is effective in both per-layer and cross-layer (shared dictionary) compression scenarios, consistently outperforming regular SVD, ASVD Yuan et al. (2023), SVD-LLM Wang et al. (2025a), and Basis Sharing Wang et al. (2024) strategies as well as state-of-the-art structure-pruning methods Ma et al. (2023); Shopkhoev et al. (2025).
2 Related Work
Research on Transformer compression covers pruning, quantization, distillation, and matrix factorization. Below we provide a brief overview of the recent progress in these directions.
Early work on pruning showed substantial redundancy in deep nets Han et al. (2015); Frankle & Carbin (2019); Sanh et al. (2020). For LLMs, SparseGPT performs one-shot post-training pruning and achieves high sparsity with small quality drop Frantar & Alistarh (2023). Wanda reduces overhead further with simple activation-based rules Sun et al. (2023). LLM-Pruner removes blocks using gradient-guided criteria and recovers accuracy with brief adapter tuning Ma et al. (2023), while ReplaceMe substitutes multiple transformer blocks with a single linear transformation Shopkhoev et al. (2025).
Quantization reduces precision to reduce memory consumption and accelerate inference. For weight-only post-training quantization, LLM.int8 enables INT8 inference via vector-wise quantization with a path for outliers Dettmers et al. (2022); GPTQ/OPTQ minimize output error using Hessian to reach 3–4 bit weights Frantar et al. (2022; 2023); AWQ derives scales from activations and protects a small set of salient weights Lin et al. (2024), and SpQR stores outliers at higher precision to maintain quality at 3–4 bits Dettmers et al. (2024). For quantizing both weights and activations, SmoothQuant rebalances channel ranges to achieve accurate W8A8 across matrix multiplications Xiao et al. (2023), while QuaRot applies orthogonal rotations to suppress outliers, enabling end-to-end 4-bit inference including the KV cache Ashkboos et al. (2024).
The general framework on knowledge distillation by Hinton et al. and sequence-level distillation by Kim and Rush established strong baselines Hinton et al. (2015); Kim & Rush (2016). DistilBERT demonstrates task-agnostic compression for Transformers Sanh et al. (2019); TinyBERT uses a two-stage distillation procedure Jiao et al. (2020); MobileBERT introduces a teacher-guided compact architecture Sun et al. (2020); Patient-KD leverages multi-layer hints Sun et al. (2019); MiniLM and MiniLMv2 distill self-attention relations Wang et al. (2020; 2021). BERT-of-Theseus compresses models via progressive module replacement Xu et al. (2020). Orca and Orca 2 show that richer teacher signals such as explanations can improve reasoning in smaller students Mukherjee et al. (2023); Mitra et al. (2023).
Classic low-rank approximations reduce parameters and FLOPs with limited loss Denton et al. (2014). DRONE considers an objective related to the output activation for the weight approximation, Chen et al. (2021). FWSVD introduces Fisher-weighted reconstruction to emphasize important directions Hsu et al. (2022), while ASVD adapts truncation using activation-aware transforms and layer sensitivity Yuan et al. (2023). SVD-LLM utilizes truncation-aware whitening , while SVD-LLM v2 optimizes budget allocation across layers Wang et al. (2025a; b). Basis Sharing reuses one part of low-rank factorization across layers and learns layer-specific coefficients to exploit inter-layer redundancy Wang et al. (2024).
Dictionary learning and sparse coding. Dictionary learning factorizes original weight into dictionary and sparse codes, yielding a union-of-subspaces model with strong results in vision and image compression Engan et al. (1999); Aharon et al. (2006); Mairal et al. (2010); Gregor & LeCun (2010); Elad (2010). In NLP, GroupReduce explores block-wise low-rank/dictionary-style compression Chen et al. (2018), and tensorized Transformers factorize attention and embeddings Ma et al. (2019). Recent work also targets the KV-cache by learning universal dictionaries and decoding with OMP during inference Kim et al. (2024). Complementary to these directions, a recent work proposes cross-layer weight sharing via a matrix-based dictionary learning formulation for Transformer attention, directly exploiting inter-layer redundancy Zhussip et al. (2025). Cross-layer/shared-basis schemes in attention are thus closely related Wang et al. (2024). A comprehensive, training-free approach to weight-space dictionary learning for LLMs — at both per-layer and cross-layer levels — remains underexplored and is the focus of this work.
Figure 1: Left side: weight factorization methods using low-rank decomposition. Low-rank approximation decomposes a matrix into two dense matrices of lower rank. Right side: proposed CoSpaDi. A dictionary of $k$ atoms and a column-sparse coefficient matrix are employed. No restrictions on size of $k$ (undercomplete : $k<d_{1}$, complete: $k=d_{1}$ or overcomplete : $k>d_{1}$ dictionaries are possible), while sparsity is defined by $s$ non-zero elements per column of the coefficient matrix.
3 Proposed Method
In this section, we first provide a conceptual link between low-rank weight approximation and dictionary learning. Then, we introduce the proposed Sparse Dictionary Learning (SDL) strategy for LLM compression covering all practical aspects including data-aware SDL, dictionary-sharing among layers, compression ratios and inference complexity.
3.1 Low-Rank Weight Approximation
A widely adopted strategy to reduce the parameter count in LLMs is to approximate weight matrices $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ — representing the parameters of network layers — with matrices of reduced rank $r<\min(d_{1},d_{2})$. Such a low-rank approximation can be derived as follows:
$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\operatorname*{rank}\left(\tilde{\mathbf{W}}\right)=r}\left\|\mathbf{W}-\tilde{\mathbf{W}}\right\|_{F}.$ (1)
According to the Eckart–Young–Mirsky theorem Eckart & Young (1936), the orthogonal projection to the space of $r$-rank matrices admits an analytical solution. Specifically, if $\mathbf{W}$ admits the singular value decomposition $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, with $\mathbf{U}\in\mathbb{R}^{d_{1}\times k}$, $\bm{\Sigma}\in\mathbb{R}^{k\times k}$ and $\mathbf{V}\in\mathbb{R}^{d_{2}\times k}$ with $k=\min\left(d_{1},d_{2}\right)$, then the minimizer of Eq. (1) can be expressed as: $\tilde{\mathbf{W}}^{\star}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$, where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ contains the first $r$ left singular vectors, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ holds the top-$r$ singular values, and $\mathbf{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ contains the first $r$ right singular vectors of $\mathbf{W}$, respectively.
There is a deep connection between this problem and the principal component analysis (PCA) Bishop & Nasrabadi (2006). Specifically, consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$. We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$. The optimal basis and coefficients can be found by minimizing the total approximation error:
$\mathcal{J}=\sum\limits_{j=1}^{d_{2}}\left\|\mathbf{w}_{j}-\sum\limits_{i=1}^{r}c_{i,j}\bm{b}_{i}\right\|_{2}^{2}\equiv\left\|\mathbf{W}-\mathbf{B}\mathbf{C}\right\|_{F}^{2},$ (2)
subject to the orthogonality constraint $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}$, where $\mathbf{B}=\begin{bmatrix}\bm{b}_{1},\ldots,\bm{b}_{r}\end{bmatrix}\in\mathbb{R}^{d_{1}\times r}$ is the basis matrix, $\mathbf{C}\in\mathbb{R}^{r\times d_{2}}$ is the coefficient matrix with entries $c_{i,j}$, and $\mathbf{I}\in\mathbb{R}^{r\times r}$ is the identity matrix.
The optimal pair $\left(\mathbf{B}^{*},\mathbf{C}^{*}\right)$ is obtained by solving:
$\mathbf{B}^{*},\mathbf{C}^{*}=\operatorname*{arg\,min}_{\mathbf{B},\mathbf{C}}\mathcal{J}\,\,\mbox{s.t}\,\,\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}.$ (3)
The solution is given by $\mathbf{B}^{*}=\mathbf{U}_{r}$ and $\mathbf{C}^{*}=\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$ (we refer to Appendix A.1 for the proof). This result highlights that in the low-rank factorization $\mathbf{W}\approx\mathbf{B}\mathbf{C}$, the matrix $\mathbf{B}$ corresponds to the basis of the shared $r$-dimensional subspace in which the columns of $\mathbf{W}$ approximately lie, and $\mathbf{C}$ contains their coordinate vectors (coefficients) with respect to that basis.
3.2 Sparse Dictionary Learning
Motivated by this interpretation, we propose an alternative strategy for compressing the weights $\mathbf{W}$, based on the sparse dictionary learning methodology. Specifically, rather than modeling each column of the weight matrix as a linear combination of basis vectors $\bm{b}_{i}$, we aim to learn a dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and approximate each column $\mathbf{w}_{j}$ as a linear combination of dictionary atoms $\mathbf{d}_{i}\in\mathbb{R}^{d_{1}}\text{ with }i=1,\ldots,k$.
Although the two strategies appear similar, they differ in two key respects. First, in dictionary learning, unlike low-rank approximation, only a subset of atoms is used to represent each weight vector $\mathbf{w}_{j}$. This principle, known in the literature as sparse coding Lee et al. (2006), is motivated by the observation that learned dictionaries typically yield sparse representations: each signal (column) activates only a few atoms. In other words, instead of approximating the weight matrix $\mathbf{W}$ as the product of two dense matrices, the orthogonal basis $\mathbf{B}$ and the coefficient matrix $\mathbf{C}$, we approximate it as the product of a dense dictionary matrix $\mathbf{D}$ and a column-wise sparse coefficient matrix $\mathbf{S}$, where each column $\mathbf{s}_{j}$ indicates which atoms contribute to the representation of $\mathbf{w}_{j}$, and with what weights. Second, and of significant practical consequence, we do not impose orthogonality constraints on the dictionary atoms. This grants greater representational flexibility, allowing the model to adapt more effectively to the intrinsic structure of the weight matrix.
Finally, under the proposed strategy, rather than enforcing a shared low-dimensional subspace for all columns, each weight vector $\mathbf{w}_{j}$ is represented within its own low-dimensional subspace spanned by the atoms it activates. Consequently, the entire matrix $\mathbf{W}$ is modeled as a union of low-dimensional subspaces, introducing additional representational capacity and enabling lower approximation error compared to a single shared subspace.
Formally, the dictionary learning problem can be expressed as:
$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{W}-\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2},$ (4)
where $\mathbf{s}_{j}$ denotes the $j$-th column of the sparse coefficient matrix $\mathbf{S}$, and $\left\|\cdot\right\|_{0}$ denotes the $\ell_{0}$ pseudo-norm, which counts the number of nonzero entries in a vector.
The optimization problem in Eq. (4) is NP-hard in its original form and admits no closed-form solution. Nevertheless, it has been extensively studied, and several efficient algorithms have been proposed to obtain high-quality approximate solutions. Among these, K-SVD Aharon et al. (2006) is a well-established algorithm widely adopted across diverse applications.
K-SVD alternates between two main steps: (a) Sparse coding: Each signal $\mathbf{w}_{j}$ is approximated as a sparse linear combination of dictionary atoms (typically via orthogonal matching pursuit or LASSO). (b) Dictionary update: Each atom $\mathbf{d}_{i}$ is refined to better fit the data while preserving the sparsity pattern in $\mathbf{S}$. This is performed sequentially per atom: the algorithm identifies the subset of signals that activate $\mathbf{d}_{i}$, computes the residual error over those signals, and applies a rank-1 SVD to jointly update the atom and its associated coefficients.
3.3 CoSpadi: Activation-Aware SDL
In several recent works, including Chen et al. (2021), it has been observed that LLM weight matrices generally do not exhibit intrinsic low-rank structure. Consequently, direct low-rank approximation of the weights often leads to significant performance degradation. However, under the hypothesis that the input latent features to a given layer reside in a low-dimensional subspace, it remains feasible and often effective to approximate the corresponding layer weights as low-rank. This approach is well-motivated: the goal is not to preserve the weights in isolation, but rather to maintain the fidelity of the activations they produce when applied to low-dimensional inputs.
Within this framework, an effective approximation of $\mathbf{W}$ can be obtained by minimizing the output error:
$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\right\|_{F}\quad\text{s.t.}\quad\operatorname*{rank}(\mathbf{X}\tilde{\mathbf{W}})=r,$ (5)
where $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ is a matrix of $N$ calibration input vectors $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d_{1}}\right\}_{i=1}^{N}$. This minimization admits the analytical solution: $\tilde{\mathbf{W}}=\mathbf{L}^{-1}\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}=\mathbf{B}\mathbf{C}$, where $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is a non-singular transformation matrix derived from the calibration data (see Appendix A.2 for the derivation).
Motivated by this data-aware perspective, we propose adapting our sparse dictionary learning framework accordingly. Rather than directly approximating $\mathbf{W}$ as $\tilde{\mathbf{W}}=\mathbf{D}\mathbf{S}$, we instead minimize the reconstruction error of the output activations $\mathbf{Z}=\mathbf{X}\mathbf{W}$:
$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (6)
To simplify this optimization, let $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}\in\mathbb{R}^{N\times d_{1}}$ be a column-orthogonal matrix, i.e., $\mathbf{Y}^{\mathsf{T}}\mathbf{Y}=\mathbf{I}_{d_{1}}$, obtained by applying a linear transformation $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ to $\mathbf{X}$. We assume $N\geq d_{1}$ and that $\mathbf{X}$ has full column rank — conditions typically satisfied with sufficient calibration data. The matrix $\mathbf{L}$ can be computed via QR decomposition, SVD of $\mathbf{X}$, or Cholesky/eigen-decomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
Introducing the auxiliary variables $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}$ and $\mathbf{D}_{L}=\mathbf{L}\mathbf{D}$, and noting that the Frobenius norm is invariant under left-multiplication by a column-orthogonal matrix, we can reformulate Eq. (6) as:
$\mathbf{D}_{L}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D}_{L},\mathbf{S}}\left\|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (7)
Since $\mathbf{L}$ is invertible, the final structured approximation of the original weights is given by:
$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S},\quad\text{where}\quad\mathbf{D}_{a}=\mathbf{L}^{-1}\mathbf{D}_{L}$ (8)
is the activation-aware learned dictionary. Pseudocode for the full procedure is provided in Appendix A.3.
Cross-Layer SDL
Although most compression techniques focus on intra-layer optimizations, the repetitive layered structure of LLMs suggests the presence of significant inter-layer redundancy. To exploit this, we propose sharing a single dictionary across weight matrices of the same type from multiple layers. Specifically, let $\mathcal{G}$ denote a group of layer indices (e.g., all attention projection layers or all FFN layers). We form a grouped weight matrix by horizontally concatenating the corresponding layer weights: $\mathbf{W}_{G}=\begin{bmatrix}\mathbf{W}_{\ell_{1}},&\mathbf{W}_{\ell_{2}},&\ldots,&\mathbf{W}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{d_{1}\times(d_{2}\cdot L)},$ where each $\mathbf{W}_{\ell}\in\mathbb{R}^{d_{1}\times d_{2}}$, and $L=|\mathcal{G}|$ is the number of layers in the group. We then apply the same activation-aware dictionary learning procedure as before, but now to $\mathbf{W}_{G}$. To ensure the approximation remains data-aware across all layers in the group, we construct a grouped calibration matrix by vertically stacking the corresponding input batches: $\mathbf{X}_{G}=\begin{bmatrix}\mathbf{X}_{\ell_{1}};&\mathbf{X}_{\ell_{2}};&\ldots;&\mathbf{X}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{(N\cdot L)\times d_{1}}$ where each $\mathbf{X}_{\ell}\in\mathbb{R}^{N\times d_{1}}$ is the calibration input for layer $\ell$. The transformation matrix $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is then computed from $\mathbf{X}_{G}$ (e.g., via QR or Cholesky decomposition), ensuring that $\mathbf{Y}_{G}=\mathbf{X}_{G}\mathbf{L}^{-1}$ is column-orthogonal. The optimization proceeds by solving Eq. (7) with $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}_{G}$, yielding a shared dictionary $\mathbf{D}_{a}$ and a sparse coefficient matrix $\mathbf{S}_{G}\in\mathbb{R}^{k\times(d_{2}\cdot L)}$.
This approach, inspired by Wang et al. (2024), reduces memory overhead by amortizing the dictionary cost across multiple layers, while preserving activation fidelity through data-aware calibration. Each layer’s compressed weights are then recovered by slicing the corresponding block from $\tilde{\mathbf{W}}_{G}=\mathbf{D}_{a}\mathbf{S}_{G}$.
Compression ratio.
For the low-rank setting, the compression ratio is $\gamma^{\mathrm{LR}}\coloneqq 1-\tfrac{r(d_{1}+d_{2})}{d_{1}d_{2}}$, where $r$ is the retained rank. For CoSpaDi, we store a dictionary and sparse codes plus a binary mask (details are provided in section A.4), giving
$\gamma^{\mathrm{SD}}\coloneqq 1-\tfrac{\;\overbrace{d_{1}k}^{\text{dict. (bf16)}}+\overbrace{sd_{2}}^{\text{codes (bf16)}}+\overbrace{(kd_{2})/16}^{\text{mask (1 bit/entry)}}\;}{d_{1}d_{2}}.$ (9)
Unlike the low-rank case, $\gamma^{\mathrm{SD}}$ depends on two knobs — the number of atoms $k$ and sparsity $s$ — allowing us to trade off model capacity and storage at fixed budget. We parameterize this trade-off by the ratio $\rho\coloneqq k/s$, which uniquely determines $(k,s)$ at a target $\gamma^{\mathrm{SD}}$:
$k=\frac{(1-\gamma^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}+\tfrac{d_{2}}{16}\,},\qquad s=\tfrac{k}{\rho}.$ (10)
Inference Complexity
In terms of computational efficiency, low-rank and dictionary learning exhibit distinct inference-time complexity profiles: the former has a cost of $Tr(d_{1}+d_{2})$, whereas the latter – when exploiting sparsity and reusing inner products – achieves $Td_{1}K_{\text{active}}+Tsd_{2}$ (see section A.5), potentially yielding superior efficiency under favorable sparsity patterns. Although both methods share identical theoretical complexity under matched compression ratios, practical inference latency varies significantly due to factors such as the number of active atoms, indexing overhead, and hardware-specific kernel efficiency. Further details regarding complexity derivations are provided in Appendix A.5.
4 Experiments
In this section, we present our experimental setup. We first compare CoSpaDi with low-rank baselines in the per-layer setting, where each linear weight matrix is compressed independently. We then investigate cross-layer sharing, analyze the impact of coefficient matrix quantization, and conduct ablation studies on the $ks$-ratio.
4.1 Experimental Setup
We evaluate our method in both per-layer and cross-layer settings. For per-layer evaluations, we consider LLaMA-3.2 1B, Qwen-3 0.6B, LLaMA-3 8B, and Qwen-3 8B. For the cross-layer study, we follow the Basis Sharing Wang et al. (2024) setup and conduct experiments on LLaMA-2 7B. All models are evaluated in a zero-shot setting on PIQA Bisk et al. (2019), HellaSwag Zellers et al. (2019), OpenAI LAMBADA Paperno et al. (2016), ARC-easy and ARC-challenge Clark et al. (2018), SciQ Welbl et al. (2017), Race Lai et al. (2017), and MMLU Hendrycks et al. (2021), while perplexity is additionally reported on WikiText Merity et al. (2016) and LAMBADA-OpenAI. Compression is applied with target ratios of 0.2-0.5 with 0.1 step. For methods requiring calibration data, we randomly sample 256 sequences from the RefinedWeb dataset Penedo et al. (2023), each consisting of 1024 tokens. Unless specified otherwise, we compress all dense linear projections in self-attention (Q/K/V/O) and feed-forward network (gate/up/down). Embeddings and lm_head are left intact.
For low-rank methods the rank is uniquely defined with $\gamma^{LR}$ and we floored it to the nearest integer. For CoSpaDi we fixed $\rho=2$ ($k/s$ ratio) for all experiments, so $k$ and $s$ were obtained according to Eq. (10) and then were floored to the nearest integers.
In CoSpaDi we employ K-SVD using power iterations instead of full-svd for the dictionary updates. Specifically, we use 60 K-SVD iterations and 8 power iterations to ensure stable convergence. Further implementation details and convergence analysis are provided in Appendix A.8.
4.2 Ablation Study
Influence of $k/s$-ratio.
In the proposed CoSpaDi we can redistribute capacity across both $k$ and $s$ while preserving predefined compression ratio by varying the $\rho$ value ($k/s$). We verified its influence on the Llama 3.2 1B and report the metrics in Fig. 2.
Figure 2: Dual-axis plot showing average accuracy (  solid lines, left axis) and perplexity (--- dashed lines, right axis, logarithmic scale with inverted direction) as functions of $\rho$ for Llama3.2-1B under three compression levels: 0.2, 0.3 and 0.4. Perplexity decreases upward due to axis inversion.
CR Bitwidth Avg. Acc. PPL
0.1686 bFP16 0.6198 1.94E+01
0.1843 bFP15 0.6195 1.95E+01
0.2001 bFP14 0.6176 1.97E+01
0.2726 bFP16 0.5408 4.34E+01
0.2864 bFP15 0.5394 4.37E+01
0.3002 bFP14 0.5373 4.46E+01
0.3765 bFP16 0.4096 1.80E+02
0.3883 bFP15 0.4086 1.82E+02
0.4002 bFP14 0.4069 1.84E+02
0.4804 bFP16 0.2846 8.23E+02
0.4902 bFP15 0.2838 8.15E+02
0.5001 bFP14 0.2837 8.10E+02
Table 1: Results on truncation of bfloat16 mantissa bits of coefficient matrix with reported average accuracy (Avg. Acc.) and Wiki Text word perplexity (PPL) for Llama3-8B resulting in different compression ratios (CR).
From the provided plots we can observe that depending on the compression ratio the optimal $ks$-ratio differs, thus, for simplicity in further experiments we select in all cases $\rho=2$.
Data-Free and Data-Aware Scenarios We argue that our proposed CoSpaDi allows for a more flexible representation compared to low-rank approximation, regardless of whether data-free or data-aware scenarios are considered. To prove this claim we performed an ablation study on Llama3-1B for different compression levels. We selected SVD for the data-free scenario and SVD-LLM for the data-aware case, while varying the compression ratio from 0.2 to 0.5. In Table 4.2 we report the results only for 0.2-0.5 compression. These results clearly indicate that in both data-free and data-aware settings the SDL-based methods outperform low-rank baselines by a wide margin, while CoSpaDi outperforms all methods.
Table 2: SDL-based methods comparison vs low-rank counterparts in data-free and data-aware scenarios on Llama3.2-1B at different compression ratios (CR). We denote $\text{CoSpaDi}^{\dagger}$ as the proposed method without using calibration data. Best results are provided in bold.
[…]
Method Data-Aware CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
[…]
SVD x 0.5201 0.2566 0.0003 0.2471 0.2210 0.2000 0.2144 0.2537 0.2391 2.93E+06 4.56E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5174 0.2635 0.0002 0.2538 0.2534 0.2100 0.2182 0.2413 0.2447 3.33E+05 2.16E+06
SVD-LLM v 0.6213 0.3643 0.2443 0.3603 0.2509 0.6490 0.2900 0.2298 0.3762 1.69E+02 1.70E+02
CoSpaDi v 0.2 0.6605 0.4288 0.3839 0.3994 0.2602 0.7160 0.3167 0.2483 0.4267 6.37E+01 3.51E+01
SVD x 0.5234 0.2564 0.0002 0.2416 0.2346 0.1950 0.2191 0.2698 0.2425 1.09E+06 3.92E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5049 0.2626 0.0002 0.2449 0.2611 0.2180 0.2153 0.2552 0.2453 2.06E+05 4.34E+06
SVD-LLM v 0.5565 0.3006 0.0910 0.3047 0.2150 0.4590 0.2583 0.2323 0.3022 5.90E+02 2.47E+03
CoSpaDi v 0.3 0.5691 0.3241 0.1822 0.3194 0.2210 0.5670 0.2804 0.2308 0.3368 2.89E+02 6.59E+02
SVD x 0.5277 0.2590 0.0001 0.2386 0.2133 0.1990 0.2220 0.2693 0.2411 1.23E+06 4.17E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5103 0.2626 0.0001 0.2546 0.2688 0.2130 0.2115 0.2549 0.2470 3.05E+06 3.67E+07
SVD-LLM v 0.5180 0.2725 0.0126 0.2685 0.2287 0.3230 0.2440 0.2297 0.2621 1.58E+03 3.30E+04
CoSpaDi v 0.4 0.5348 0.2824 0.0380 0.2778 0.2295 0.3690 0.2402 0.2310 0.2753 7.96E+02 9.23E+03
SVD x 0.5392 0.2566 0.0001 0.2441 0.2039 0.2040 0.2077 0.2690 0.2406 1.22E+06 1.88E+07
$\text{CoSpaDi}^{\dagger}$ x 0.5120 0.2607 0.0001 0.2572 0.2654 0.2280 0.2211 0.2704 0.2519 6.09E+05 1.08E+07
SVD-LLM v 0.5109 0.2663 0.0004 0.2609 0.2594 0.2610 0.2392 0.2303 0.2536 3.13E+03 1.03E+05
CoSpaDi v 0.5 0.5169 0.2703 0.0031 0.2626 0.2398 0.2950 0.2421 0.2326 0.2578 1.80E+03 7.26E+04
Quantization of Sparse Coefficient Matrix We investigated the quantization of the coefficient matrix $\mathbf{S}$ in the post-training regime to leverage the memory add-on due to the requirement of storing the indices of the nonzero values. We utilized a naive mantissa bit truncation of the original bfloat16 coefficient values. In Table 1 we report how truncation affects the performance at different compression levels.
According to the obtained results, truncation of 2 bits leads to negligible drop. Thus, in all our experiments CoSpaDi consists of bfloat16 dictionary and coefficient matrices but is evaluated with the truncated version of the sparse coefficient matrix using 14-bits.
4.3 Main Results
Per-Layer Scenario:
In this section we apply CoSpaDi for each projection layer independently and compare it against current sota training-free low-rank method – SVD-LLM Wang et al. (2025a).
(a) Accuracy & Perplexity for LLaMA3.2-1B
(b) Accuracy & Perplexity for Qwen3-0.6B
Figure 3: Average benchmark accuracy and WikiText perplexity for (a) LLaMA-3.2-1B and (b) Qwen-3 0.6B using SVD-LLM and CoSpaDiwith respect to compression ratio.
From the provided plots CoSpaDi significantly outperforms SVD-LLM in both avg. accuracy and perplexity for small Llama3 and Qwen3 models. We further investigate how our method scales to larger models of the same families with 8B parameters. The related results are provided in Table 4.3.
We have also performed comparisons on the same models and compression levels against other training-free compression strategies. In particular we compare CoSpaDi against ASVD Yuan et al. (2023), ReplaceMe Shopkhoev et al. (2025) and LLMPruner Ma et al. (2023), with the last two being sota pruning methods. These results are provided in Appendix A.6
Table 3: Performance comparison of CoSpaDi vs sota SVD-LLM on Llama3-8B and Qwen3-8B at different compression levels on different benchmarks. Best results are highlighted with bold.
[…]
Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
[…]
SVD-LLM 0.7106 0.5837 0.5925 0.5551 0.3396 0.8640 0.3550 0.3262 0.5408 4.07E+01 1.09E+01
CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
SVD-LLM 0.6578 0.4640 0.3806 0.4188 0.2765 0.7000 0.3177 0.2715 0.4358 1.47E+02 6.09E+01
CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
SVD-LLM 0.6028 0.3450 0.1143 0.3237 0.2449 0.4420 0.2574 0.2305 0.3201 5.49E+02 1.30E+03
CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
SVD-LLM 0.5381 0.2790 0.0041 0.2731 0.2389 0.2660 0.2230 0.2297 0.2565 1.37E+03 2.80E+04
CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
[…]
SVD-LLM 0.7383 0.6391 0.6218 0.6869 0.4573 0.9010 0.4048 0.5473 0.6246 2.06E+01 6.40E+00
CoSpaDi 0.2 0.7650 0.6804 0.6559 0.7218 0.4889 0.9320 0.4067 0.6075 0.6573 1.81E+01 4.88E+00
SVD-LLM 0.7035 0.5522 0.5377 0.5926 0.3712 0.8720 0.3837 0.4482 0.5576 2.74E+01 1.07E+01
CoSpaDi 0.3 0.7242 0.6050 0.6255 0.6385 0.4121 0.8840 0.3952 0.5127 0.5997 2.29E+01 6.29E+00
SVD-LLM 0.6627 0.4458 0.3790 0.4503 0.2807 0.7730 0.3531 0.2901 0.4544 4.30E+01 3.59E+01
CoSpaDi 0.4 0.6888 0.4898 0.4991 0.4941 0.2986 0.8200 0.3675 0.3660 0.5030 3.59E+01 1.47E+01
SVD-LLM 0.6023 0.3543 0.1956 0.3363 0.2278 0.6670 0.2909 0.2305 0.3631 8.72E+01 2.69E+02
CoSpaDi 0.5 0.6159 0.3798 0.2791 0.3565 0.2295 0.6860 0.3167 0.2305 0.3868 6.82E+01 1.10E+02
Cross-Layer Scenario
In this section, we validate CoSpaDi on the cross-layer compression scenario, where a common dictionary is shared across different layers. To ensure a fair comparison, we adopt the identical layer selection and grouping procedure introduced in Basis Sharing Wang et al. (2024) (for a detailed description we refer to Appendix A.7) and compare the performance of the two methods on the Llama2-7B model Touvron et al. (2023b). As shown in Table LABEL:tab:main_results_group, CoSpaDi consistently outperforms Basis Sharing by a large margin across all benchmarks and compression rates (0.2–0.5). This performance gap suggests that CoSpaDi’s sparsity-aware, dictionary-based decomposition better preserves task-critical features under aggressive compression.
We believe that a more sophisticated layer grouping strategy, based on feature map similarity or end-to-end differentiable search, could work better with CoSpaDi and further widen this performance margin, but we leave this as a future research direction. We provide details for $k$ (dictionary size) and $s$ (sparsity) parameters in the Appendix A.7.
Appendix A Appendix
A.1 Derivation of the optimal pair of basis and coefficient matrices
We are seeking the optimal pair $\left(\mathbf{B}^{\star},\mathbf{C}^{\star}\right)$ that minimizes the constrained problem in Eq. (3). First, we rewrite the objective $\mathcal{J}$ in its equivalent form:
$\mathcal{J}=\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}\right)-2\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{B}\mathbf{C}\right)+\operatorname*{tr}\left(\mathbf{C}^{\mathsf{T}}\mathbf{B}^{\mathsf{T}}\mathbf{B}\mathbf{C}\right).$ (11)
Next, we consider the basis matrix $\mathbf{B}$ as fixed and compute the gradient of the objective w.r.t the matrix coefficient $\mathbf{C}$ as
$\nabla_{\mathbf{C}}\mathcal{J}=2\mathbf{B}^{\mathsf{T}}\mathbf{B}\mathbf{C}-2\mathbf{B}^{\mathsf{T}}\mathbf{W}.$ (12)
Setting the gradient to zero and taking into account the constraint $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}$, we can recover the optimum matrix coefficient as: $\mathbf{C}=\mathbf{B}^{\mathsf{T}}\mathbf{W}$. Now, putting back $\mathbf{C}$ in Eq. (11) we get:
$\mathcal{J}$ $=\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}\right)-2\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{B}\mathbf{B}^{\mathsf{T}}\mathbf{W}\right)+\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{B}\mathbf{B}^{\mathsf{T}}\mathbf{B}\mathbf{B}^{\mathsf{T}}\mathbf{W}\right)$
$=\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}\right)-\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{B}\mathbf{B}^{\mathsf{T}}\mathbf{W}\right)\,\,\mbox{(from the constraint $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}$)}$
$=\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}\right)-\operatorname*{tr}\left(\mathbf{B}^{\mathsf{T}}\mathbf{W}\mathbf{W}^{\mathsf{T}}\mathbf{B}\right).$ (13)
Based on this we can recover the optimum basis matrix $\mathbf{B}$ as the maximizer of the constrained problem:
$\mathbf{B}^{\star}=\operatorname*{arg\,max}_{\mathbf{B}}\operatorname*{tr}\left(\mathbf{B}^{\mathsf{T}}\mathbf{W}\mathbf{W}^{\mathsf{T}}\mathbf{B}\right)\,\,\mbox{s.t}\,\,\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}.$ (14)
The above maximization problem enjoys a closed-form solution Fan (1949), which is fully defined by the eigenvalues of the matrix $\mathbf{P}=\mathbf{W}\mathbf{W}^{\mathsf{T}}$. Specifically, the matrix $\mathbf{P}\in\mathbb{R}^{d_{1}\times d_{1}}$, which is symmetric and positive semi-definite, admits the eigenvalue decomposition $\mathbf{P}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\mathsf{T}}$, with $\mathbf{U}\in\mathbb{R}^{d_{1}\times d_{1}}$ holding the eigenvectors of $\mathbf{P}$ in its columns. Then, the maximizer of Eq. (14) is recovered as $\mathbf{B}^{\star}=\mathbf{U}_{r}$ where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ is a reduced version of $\mathbf{U}$ formed with the $r$ eigenvectors corresponding to the largest eigenvalues of $\mathbf{P}$. A useful observation is that the eigenvectors of $\mathbf{P}$ exactly match the left-singular vectors of $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$. Indeed, if $\mathbf{W}$ admits the singular value decomposition $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, then we have that: $\mathbf{P}=\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{U}\bm{\Sigma}^{2}\mathbf{U}^{\mathsf{T}}\equiv\mathbf{U}\bm{\Lambda}\mathbf{U}^{\mathsf{T}}$, with $\bm{\Lambda}=\bm{\Sigma}^{2}$. Therefore, instead of performing the eigenvalue decomposition on $\mathbf{P}$ we can recover $\mathbf{U}$ and consequently $\mathbf{U}_{r}$ by computing the SVD of $\mathbf{W}$.
Finally, we can compute the optimum coefficient matrix as:
$\mathbf{C}^{\star}$ $=\left(\mathbf{B}^{\star}\right)^{\mathsf{T}}\mathbf{W}=\mathbf{U}_{r}^{\mathsf{T}}\mathbf{W}$
$=\mathbf{U}_{r}^{\mathsf{T}}\overbrace{\begin{bmatrix}\mathbf{U}_{r}&\mathbf{U}_{d-r}\end{bmatrix}}^{\mathbf{U}}\overbrace{\begin{bmatrix}\bm{\Sigma}_{r}&\bm{O}_{r\times\left(d-r\right)}\\ \mathbf{O}_{\left(d-r\right)\times r}&\bm{\Sigma}_{d-r}\end{bmatrix}}^{\bm{\Sigma}}\overbrace{\begin{bmatrix}\mathbf{V}_{r}^{\mathsf{T}}\\ \mathbf{V}_{d-r}^{\mathsf{T}}\end{bmatrix}}^{\mathbf{V}^{\mathsf{T}}}$
$=\begin{bmatrix}\mathbf{I}_{r\times r}&\mathbf{O}_{r\times\left(d-r\right)}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{r}\bm{V}_{r}^{\mathsf{T}}\\ \bm{\Sigma}_{d-r}\mathbf{V}_{d-r}^{\mathsf{T}}\end{bmatrix}$
$=\bm{\Sigma}_{r}\bm{V}_{r}^{\mathsf{T}},$ (15)
where $\bm{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ is a reduced version of $\mathbf{V}$ formed with the $r$ right singular vectors of $\mathbf{W}$ that correspond to its top-$r$ singular values, which are kept in the diagonal matrix $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$.
A.2 Data-Aware Low-Rank Weight Approximation
While the low-rank approximation of the weights $\mathbf{W}$ has been extensively used for compression tasks, in practice is not well suited to LLMs and it can lead to a severe drop of their performance . Several recent works  have suggested that instead of approximating the weights $\mathbf{W}$ with a low-rank matrix, a more efficient strategy is to model the weight activations, $\mathbf{Z}=\mathbf{X}\mathbf{W}$, as low-rank. Here, the matrix $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\ldots&\mathbf{x}_{N}\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{N\times d}$ holds in its rows the $d$-dimensional input vectors $\mathbf{x}_{n}$ with $n=1\ldots,N$, which play the role of calibration data. Under this modeling framework, we can approximate the matrix weights $\mathbf{W}$ as the minimizer of the following problem:
$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\right\|_{F}\,\mbox{s.t.}\,\operatorname*{rank}\left(\mathbf{X}\tilde{\mathbf{W}}\right)=r.$ (16)
Let us now consider $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}\in\mathbb{R}^{N\times d_{1}}$ to be a semi-orthogonal matrix (column-orthogonal matrix), that is $\mathbf{Y}^{\mathsf{T}}\mathbf{Y}=\mathbf{I}_{d_{1}}$, which is obtained by linearly transforming the matrix $\mathbf{X}$ using a non-singular matrix $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$. Here we assume that $N\geq d$ and the matrix $\mathbf{X}$ is of full rank. We note that there are different ways we can achieve this column-orthogonalization of $\mathbf{X}$. Among them we can employ the QR/SVD decomposition on $\mathbf{X}$ and the Cholesky/Eigen-value decomposition on $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ to compute a proper linear transformation $\mathbf{L}$. By using the representation $\mathbf{X}=\mathbf{Y}\mathbf{L}$ we can rewrite the problem of Eq. (16) as:
$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\mathbf{Y}\mathbf{L}\mathbf{W}-\mathbf{Y}\mathbf{L}\tilde{\mathbf{W}}\right\|_{F}\,\mbox{s.t.}\,\operatorname*{rank}\left(\mathbf{Y}\mathbf{L}\tilde{\mathbf{W}}\right)=r.$ (17)
To solve the above minimization problem we first note that due to the orthonormal columns of $\mathbf{Y}$, it can be expressed in the equivalent form:
$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\bm{\Delta}-\mathbf{L}\tilde{\mathbf{W}}\right\|_{F}\,\mbox{s.t.}\,\operatorname*{rank}\left(\mathbf{L}\tilde{\mathbf{W}}\right)=r,$ (18)
where $\bm{\Delta}=\mathbf{L}\mathbf{W}$. Next, we introduce the auxiliary matrix $\tilde{\bm{\Delta}}=\mathbf{L}\tilde{\mathbf{W}}$ and the problem in Eq. (18) becomes:
$\tilde{\bm{\Delta}}^{*}=\operatorname*{arg\,min}_{\tilde{\bm{\Delta}}}\left\|\bm{\Delta}-\tilde{\bm{\Delta}}\right\|_{F}\,\mbox{s.t.}\,\operatorname*{rank}\left(\tilde{\bm{\Delta}}\right)=r,$ (19)
which is the orthogonal projection of $\bm{\Delta}$ to the space of $r$-rank matrices. Given that $\tilde{\bm{\Delta}}^{*}=\mathbf{L}\tilde{\mathbf{W}}^{*}$ and $\mathbf{L}$ is invertible, we can now recover $\tilde{\mathbf{W}}^{*}=\mathbf{L}^{-1}\tilde{\bm{\Delta}}^{*}$.
To conclude, if $\bm{\Delta}$ admits the singular value decomposition $\bm{\Delta}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, then the optimal $r$-rank approximation of $\mathbf{W}$ that minimizes the loss in Eq. (16) can be written in the form:
$\tilde{\mathbf{W}}=\mathbf{B}\mathbf{C}=\underbrace{\mathbf{L}^{-1}\mathbf{U}_{r}}_{\mathbf{B}}\underbrace{\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}}_{\mathbf{C}}.$ (20)
We note that in this case, unlike the direct weight low-rank approximation, the matrix $\mathbf{B}=\mathbf{L}^{-1}\mathbf{U}_{r}$ does not correspond to a basis of a subspace of $\mathbb{R}^{d_{1}}$, since its columns are no longer orthonormal, that is $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{U}_{r}^{\mathsf{T}}\left(\mathbf{L}\mathbf{L}^{\mathsf{T}}\right)^{-1}\mathbf{U}_{r}\neq\mathbf{I}$ .
A.3 Pseudo Algorithm of the Proposed Method
Goal. Given a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ and a small calibration set $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$, compute an activation-aware sparse–dictionary factorization $\mathbf{W}\approx\mathbf{\widetilde{W}}=\mathbf{D}\mathbf{S}$ under a target compression ratio. The procedure consists of whitening the activation objective, alternating sparse coding and dictionary updates on the whitened weights, and a final de-whitening step.
(1) Calibration and whitening.
Compute an invertible transform $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ (e.g., via QR/SVD of $\mathbf{X}$ or Cholesky of $\mathbf{X}^{\top}\mathbf{X}$) such that $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}$ has orthonormal columns ($\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{I}$). Left-multiply $\mathbf{W}$ to obtain the whitened weights $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}$. Whitening converts the data-aware loss $\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{\widehat{W}}\|_{F}^{2}$ into a standard Frobenius objective $\|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\|_{F}^{2}$ that is amenable to dictionary learning.
(2) Initialization.
Initialize the whitened dictionary $\mathbf{D_{L}^{(0)}}\in\mathbb{R}^{d_{1}\times k}$ (e.g., random permutation of columns of $\mathbf{W}$) and set $\mathbf{S}^{(0)}=0$. The pair $(k,s)$ is set from the target compression ratio via Eq. 10 optionally using the fixed ratio $\rho=k/s$.
(3) Alternating minimization.
Repeat for $t=1,\dots,T$: (a) Sparse coding. For each column $j$, solve
$\mathbf{s_{j}^{(t)}}\;\in\;\arg\min_{\|\mathbf{s}\|_{0}\leq s}\ \big\|(\mathbf{W}_{L})_{:,j}-\mathbf{D_{L}}^{(t-1)}\mathbf{s}\big\|_{2}^{2},$
using OMP (greedy selection with orthogonal residual updates) to enforce exactly $s$ nonzeros per column. (b) Dictionary update. For each atom $i$, collect its support $\Omega_{i}=\{j:\,s_{i,j}^{(t)}\neq 0\}$ and form the residual on those columns:
$\mathbf{R}_{i}\;=\;\mathbf{W}_{L}[:,\Omega_{i}]\;-\!\!\sum_{\ell\neq i}\mathbf{D_{L,\ell}}^{(t-1)}\,\mathbf{s}_{\ell,\Omega_{i}}^{(t)}.$
Update $(\mathbf{D_{L,i}}^{(t)},\,\mathbf{s}_{i,\Omega_{i}}^{(t)})$ by the best rank-1 approximation $\mathbf{R}_{i}\approx\mathbf{u}\,\mathbf{\sigma}\,\mathbf{v}^{\top}$ (set $mD_{L,i}^{(t)}\!\leftarrow\mathbf{u}$, $\mathbf{s}_{i,\Omega_{i}}^{(t)}\!\leftarrow\mathbf{\sigma}\mathbf{v}^{\top}$). This preserves the current sparsity pattern while reducing the residual. Iterate atoms sequentially. Stop when the maximum iteration $T$ is reached or when the relative improvement falls below a tolerance.
(4) De-whitening and packing.
Map the dictionary back to activation space via $\mathbf{D_{a}}=\mathbf{L^{-1}}\mathbf{D_{L}^{(T)}}$ and set $\widetilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S}^{(T)}$. For storage, keep $\mathbf{D}_{a}$ and the $sd_{2}$ nonzero entries of $\mathbf{S}$ in bf16 along with a packed binary mask $\mathbf{M}\in\{0,1\}^{k\times d_{2}}$ for locations (one bit per entry; $\tfrac{kd_{2}}{16}$ words). This yields the compression ratio in Appendix A.4 and Eq. 9.
(5) Inference.
At runtime, apply $\widetilde{\mathbf{W}}$ as $\mathrm{matmul}(\mathbf{x},\mathbf{D}_{a}\mathbf{S})$ with sparse–dense kernels. Reuse inner products $\langle\mathbf{x},\mathbf{D_{a,:,i}}\rangle$ across columns to achieve the complexity in Appendix A.5; the number of active atoms controls the practical speedup.
(6) Cross-layer variant.
For a group of layers $\mathbf{G}$, concatenate weights horizontally $\mathbf{W_{G}}=[\mathbf{W_{\ell_{1}}}\ \cdots\ \mathbf{W_{\ell_{L}}}]$ and stack calibration batches vertically to form $\mathbf{X}_{G}$. Compute $\mathbf{L}$ from $\mathbf{X}_{G}$, run the same alternating procedure on $\mathbf{W}_{L,G}=\mathbf{L}\mathbf{W}_{G}$ to obtain a single shared $\mathbf{D}_{a}$, and slice the corresponding blocks of $\mathbf{S}_{G}$ back to per-layer coefficients.
Input :  $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$: weight matrix to compress
$\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$: calibration input data (NN samples)
$k$: dictionary size (number of atoms, $k\geq s$)
$s$: sparsity level (max nonzeros per column in $\mathbf{S}$)
$T$: number of K-SVD iterations
Output :  $\mathbf{D}_{a}\in\mathbb{R}^{d_{1}\times k}$: activation-aware dictionary
$\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$: sparse coefficient matrix
$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S}$: compressed weight matrix
Compute $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ such that $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}$ satisfies $\mathbf{Y}^{\mathsf{T}}\mathbf{Y}=\mathbf{I}_{d_{1}}$;
% e.g., via QR: $\mathbf{X}=\mathbf{Q}\mathbf{R}\Rightarrow\mathbf{L}=\mathbf{R}$
% e.g., via Cholesky: $\mathbf{X}^{\mathsf{T}}\mathbf{X}=\mathbf{C}^{\mathsf{T}}\mathbf{C}\Rightarrow\mathbf{L}=\mathbf{C}$
$\mathbf{W}_{L}\leftarrow\mathbf{L}\mathbf{W}$;
Initialize $\mathbf{D}_{L}^{(0)}\in\mathbb{R}^{d_{1}\times k}$ with random Gaussian or SVD-based atoms;
Initialize $\mathbf{S}^{(0)}\in\mathbb{R}^{k\times d_{2}}$ as zero matrix;
for $t=1$ to $T$ do
 for $j=1$ to $d_{2}$ do
    $\mathbf{s}_{j}^{(t)}\leftarrow\operatorname*{arg\,min}_{\left\|\mathbf{s}\right\|_{0}\leq s}\left\|\mathbf{W}_{L,j}-\mathbf{D}_{L}^{(t-1)}\mathbf{s}\right\|_{2}^{2}$;
    % Solve via OMP, LASSO, or thresholding
   end for
 for $i=1$ to $k$ do
    $\Omega_{i}\leftarrow\left\{j\,|\,s_{i,j}^{(t)}\neq 0\right\}$;
    if $\Omega_{i}\neq\emptyset$ then
       $\mathbf{R}_{i}\leftarrow\mathbf{W}_{L}[:,\Omega_{i}]-\sum_{l\neq i}\mathbf{d}_{L,l}^{(t-1)}\mathbf{s}_{l,\Omega_{i}}^{(t)}$;
[…]
       $\mathbf{d}_{L,i}^{(t)}\leftarrow\mathbf{u}$;
       $\mathbf{s}_{i,\Omega_{i}}^{(t)}\leftarrow\sigma\cdot\mathbf{v}^{\mathsf{T}}$;
      end if
   end for
end for
$\mathbf{D}_{a}\leftarrow\mathbf{L}^{-1}\mathbf{D}_{L}^{(T)}$;
$\tilde{\mathbf{W}}\leftarrow\mathbf{D}_{a}\mathbf{S}^{(T)}$;
return $\mathbf{D}_{a}$, $\mathbf{S}^{(T)}$, $\tilde{\mathbf{W}}$;
Algorithm 1 Pseudo algorithm of the proposed CoSpaDi which consists of two steps: (a) sparse coding to compute the coefficients and (b) sequential dictionary update step.
A.4 Derivation of the CoSpaDi Compression Ratio
We derive the expression for the compression ratio $\gamma^{\mathrm{SD}}$ of our sparse–dictionary (SD) parameterization. Let $\mathbf{W}\!\in\!\mathbb{R}^{d_{1}\times d_{2}}$ be factorized as
$\mathbf{W}\;\approx\;\mathbf{D}\mathbf{S},\qquad\mathbf{D}\in\mathbb{R}^{d_{1}\times k},\ \ \mathbf{S}\in\mathbb{R}^{k\times d_{2}},$
where each column of $\mathbf{S}$ has exactly $s$ nonzeros (column-$s$-sparse). Throughout, we store real values in bfloat16 (16 bits) as is common in modern LLMs.
Dense baseline.
A dense $\mathbf{W}$ requires $d_{1}d_{2}$ bf16 values.
Dictionary term.
The dictionary $\mathbf{D}$ stores $d_{1}k$ bf16 values.
Sparse codes.
Naively, $\mathbf{S}$ would need $kd_{2}$ values. Since $\mathbf{S}$ is column-$s$-sparse, only $sd_{2}$ values are stored. For locations, one option is COO: per nonzero we keep a row index and (redundantly) the column index. Because sparsity is fixed per column, column indices can be omitted; keeping only row indices yields $sd_{2}$ indices. With 16-bit indices, the total becomes $sd_{2}$ values ++ $sd_{2}$ indices $2sd_{2}$ 16-bit words. For typical $\rho\!\coloneqq\!k/s=2$, this equals $kd_{2}$ words—offering no savings over dense $\mathbf{S}$ storage.
Instead, we use a bit mask $\mathbf{M}\!\in\!\{0,1\}^{k\times d_{2}}$ to mark nonzero positions. This requires $kd_{2}$ bits, i.e., $\nicefrac{{kd_{2}}}{{16}}$ 16-bit words after packing. We then store $sd_{2}$ bf16 values for the nonzeros and the packed mask for their positions.
Total and ratio.
The SD parameterization thus stores
$\underbrace{d_{1}k}_{\text{dictionary}}\;+\;\underbrace{sd_{2}}_{\text{values}}\;+\;\underbrace{\tfrac{kd_{2}}{16}}_{\text{mask}}$
16-bit words. Relative to the dense baseline $d_{1}d_{2}$, the resulting compression ratio is
$\gamma^{\mathrm{SD}}\coloneqq 1-\tfrac{\;\overbrace{d_{1}k}^{\text{dict. (bf16)}}+\overbrace{sd_{2}}^{\text{codes (bf16)}}+\overbrace{(kd_{2})/16}^{\text{mask (1 bit/entry)}}\;}{d_{1}d_{2}}.$
This matches the expression used in the main text and makes explicit the dependence on the two design knobs $(k,s)$.
A.5 Low-rank and CoSpaDi Inference Complexity
Here we derive the multiplication complexity for the original weight, SVD-compressed weight, and dictionary-learning (k-SVD) compression. We count multiplications only (additions are of the same order). Let $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ be a projection matrix in some layer and $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ be an input feature map; then a dense product $\mathbf{Y}=\mathbf{X}\mathbf{W}$ costs
$O_{baseline}=Nd_{1}d_{2}.$ (21)
For low-rank, in particular, SVD compression with rank $r$ the projection matrix is approximated with two matrices $\mathbf{W}\approx\mathbf{U}\mathbf{V}$ with $\mathbf{U}\in\mathbb{R}^{d_{1}\times r}$ and $\mathbf{V}\in\mathbb{R}^{r\times d_{2}}$, resulting in the following complexity:
$O_{LR}=Nd_{1}r+Nrd_{2}=Nr(d_{1}+d_{2}).$ (22)
Sparse dictionary (SD) learning similarly represents $\mathbf{W}\approx\mathbf{D}\mathbf{S}$ with dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ of $k$ atoms and sparse coefficient matrix $\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$. Omitting sparsity of $\mathbf{S}$ will result in:
$O_{SD,dense}=Nd_{1}k+Nkd_{2}=Nk(d_{1}+d_{2}).$ (23)
Taking into account that each column $\mathbf{s}_{j}$ of $\mathbf{S}$ is $s$-sparse, the $(i,j)$ element of $\mathbf{Y}=\mathbf{X}\mathbf{D}\mathbf{S}$ is
$y_{i,j}=\sum_{k=1}^{K}\mathbf{S}_{k,j}\langle\mathbf{X}_{i,:},\mathbf{D}_{:,k}\rangle=\sum_{k\in\mathbf{S}_{j}}\mathbf{S}_{k,j}\langle\mathbf{X}_{i,:},\mathbf{D}_{:,k}\rangle,$ (24)
where $S_{j}=\mathrm{supp}(s_{j})$ and $|\mathbf{S}_{j}|=s$. The overall sparse complexity depends on whether the inner products $\langle\mathbf{X}_{i,:},\mathbf{D}_{:,k}\rangle$ are reused across columns. With the most efficient way with reuse letting $\mathbf{U}=\bigcup_{j=1}^{d_{2}}S_{j}$ and $K_{active}=|U|$ we have:
$O_{SD,sparse-reuse}=Nd_{1}K_{active}+Nsd_{2},\qquad s\leq K_{active}\leq\min(K,sd_{2}).$ (25)
With proposed truncation of 2-bits in mantissa we can omit term for storing indices of nonzero elements in $\mathbf{S}$ resulting in corrected compression ratio: ^γ^SD = 1 - d1k + s d2d1d2 The rank for the low-rank decomposition could be estimated from the compression ratio with the following equation: r=(1 - γLR) d1d2d1+ d2 For sparse dictionary based method we defined $\rho=k/s$ and, thus we can estimate both $k$ and $s$ in the following way: k=(1-^γSD) d1d2d1+d2ρ,  s=kρ. Under the same compression ratio for both SVD and CoSpaDi we obtain exactly the same complexity: O_LR=N r (d_1 + d_2) = N (1 - γ^LR) d_1 d_2 O_SD=N d_1 k + N s d_2 = N(d_1 k + d_2 kρ) = N k (d_1 + d2ρ) = N (1-^γSD) d1d2d1+d2ρ(d_1 + d2ρ)=N(1-^γ^SD) d_1d_2
$O_{LR}=O_{SD}=N(1-\gamma)d_{1}d_{2}.$ (26)
While, theoretical complexity is the same, in practice inference time depends on the sparsity structure ($K_{\text{active}}$), indexing overhead, and kernel efficiency, as can be observed from Figs. 4, 5 and   6.
Figure 4: Inference time for different projection layers of Llama3.2 1B for different compression ratios and $k/s$ ratios on A100
Figure 5: Inference time for different projection layers of Llama3 8B for different compression ratios and $k/s$ ratios on A100
Figure 6: Inference time for different projection layers of Qwen3 0.6B for different compression ratios and $k/s$ ratios on A100
A.6 Comparison with Other Training-Free Methods
We also evaluate the performance of the proposed CoSpaDi relative to other training free methods, particularly structural pruning ones and ASVD Yuan et al. (2023) which is another data-aware SVD-based method. The results are provided in Table A.6.
Across LLaMA-3 8B, our method consistently outperforms structural pruning baselines at moderate budgets. At $\text{CR}\!\approx\!0.2$, CoSpaDi attains the best average accuracy (0.618 vs. 0.580 for ReplaceMe and 0.551 for LLM-Pruner) while also delivering the lowest perplexities, indicating balanced generative and discriminative quality. At $\text{CR}\!\approx\!0.3$, CoSpaDi widens the margin in average accuracy (0.537 vs. 0.469/0.405) and avoids the severe perplexity blow-ups observed for pruning methods. Even under aggressive compression ($\text{CR}\!\approx\!0.4$–0.5), CoSpaDi remains competitive or superior on most tasks and preserves far more stable perplexity profiles, whereas ReplaceMe and LLM-Pruner exhibit task collapse (near-zero LAMBADA and extreme PPL spikes). Overall, the results support dictionary–sparsity as a more robust data-free alternative to structural pruning at matched or tighter budgets, especially when both accuracy and perplexity must be maintained.
Table 5: Comparison of the proposed CoSpaDi method with other training-free methods including ASVD Yuan et al. (2023) and state-of-the-art structured pruning methods ReplaceMe Shopkhoev et al. (2025) and LLM-Pruner Ma et al. (2023) on Llama3 8B under different compression ratios. We report accuracy on different benchmarks as well as its average and perplexity. Best results are highlighted with bold
[…]
Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
[…]
ASVD 0.2 0.5577 0.2666 0.0070 0.3237 0.2150 0.3050 0.2182 0.2347 0.2660 9.38E+04 9.89E+05
ReplaceMe 0.22 0.7307 0.6572 0.4211 0.6587 0.4369 0.8640 0.3541 0.5167 0.5799 3.37E+01 2.01E+01
LLM-Pruner 0.7546 0.6746 0.5096 0.6208 0.3660 0.8780 0.3512 0.2495 0.5505 1.60E+01 1.05E+01
CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
ASVD 0.3 0.5359 0.2595 0.0000 0.2694 0.2142 0.2190 0.2115 0.2349 0.2431 3.43E+05 1.35E+07
ReplaceMe 0.31 0.6659 0.5382 0.2401 0.5067 0.3788 0.7730 0.3397 0.3058 0.4685 6.67E+01 1.33E+02
LLM-Pruner 0.6725 0.4509 0.2090 0.4541 0.2875 0.6340 0.3014 0.2292 0.4048 3.79E+01 2.21E+02
CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
ASVD 0.4 0.5365 0.2575 0.0000 0.2504 0.2287 0.2080 0.2201 0.2431 0.2430 2.05E+05 2.57E+07
ReplaceMe 0.41 0.6170 0.4430 0.0976 0.3742 0.2747 0.6040 0.3158 0.2638 0.3738 2.30E+02 1.76E+03
LLM-Pruner 0.5033 0.2580 0.0151 0.2643 0.2577 0.2810 0.2182 0.2324 0.2537 $\infty$ 5.67E+05
CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
ASVD 0.5196 0.2547 0.0000 0.2601 0.2261 0.1980 0.2134 0.2554 0.2409 1.72E+06 1.72E+07
ReplaceMe 0.5724 0.3635 0.0367 0.2976 0.2543 0.3810 0.2785 0.2300 0.3017 6.99E+02 5.03E+04
LLM-Pruner 0.5038 0.2623 0.0043 0.2630 0.2637 0.2560 0.2172 0.2371 0.2509 1.24E+03 1.23E+06
CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
A.7 Cross-Layer Scenario
In Table 6, we report the dictionary size ($k$), sparsity level ($s$), and input/output weight dimensions for each linear transform under a fixed $k/s$ ratio. This includes the query, key, value, and output projections within the attention block, as well as the up, down, and gated projections in the gated MLP block. The dictionary and sparsity dimensions are adjusted according to the target compression rate.
Table 6: Compression configurations for Llama2 7B weight matrices across varying compression rates (20%–50%). Each row specifies sparsity parameters ($k$, $s$, $k/s$ ratio) and weight dimensions ($d_{1}$, $d_{2}$) for different weight types (Query, Key, Value, Up, Gate, Down, Out), grouped by compression rate and group size (1 or 2).
Weight Type Compression Rate Group size Dictionary, $k$ Sparsity, $s$ $k/s$ ratio d1 d2
Query 20% 2 3276 1638 2 4096 8192
Key 3276 1638 2 4096 8192
Value 3276 1638 2 4096 8192
Up 4776 2388 2 4096 22016
Gate 4776 2388 2 4096 22016
Down 1 2762 1381 2 11008 4096
Out 2184 1092 2 4096 4096
Query 30% 2 2866 1433 2 4096 8192
Key 2866 1433 2 4096 8192
Value 2866 1433 2 4096 8192
Up 4178 2089 2 4096 22016
Gate 4178 2089 2 4096 22016
Down 1 2416 1208 2 11008 4096
Out 1910 955 2 4096 4096
Query 40% 2 2456 1228 2 4096 8192
Key 2456 1228 2 4096 8192
Value 2456 1228 2 4096 8192
Up 3582 1791 2 4096 22016
Gate 3582 1791 2 4096 22016
Down 1 2072 1036 2 11008 4096
Out 1638 819 2 4096 4096
Query 50% 2 2048 1024 2 4096 8192
Key 2048 1024 2 4096 8192
Value 2048 1024 2 4096 8192
Up 2984 1492 2 4096 22016
Gate 2984 1492 2 4096 22016
Down 1 1726 863 2 11008 4096
Out 1364 682 2 4096 4096
A.8 K-SVD and Power Iteration Analysis
We conducted ablation studies to assess the effect of the number of K-SVD iterations and power iterations on performance using Llama3.2-1B with fixed $\rho=2$. The left plot in Fig. 7 shows that average accuracy stabilizes after roughly 50 K-SVD iterations, while perplexity continues to decrease slightly before flattening out. The right plot of Fig. 7 indicates that very few power iterations are sufficient for stable convergence: performance improves sharply up to around 5 iterations, after which additional iterations yield minimal benefit. Based on these results, we fixed the number of K-SVD iterations to 60 and power iterations to 8 in our final experiments, which provides a good balance between accuracy, perplexity, and computational efficiency.
Figure 7: Average benchmark accuracy and WikiText perplexity with respect to the number of K-SVD iterations (left) and the number of power iterations (right) for Llama3.2-1B with $\rho=2$
A.9 Use of Large Language Models
The authors acknowledge the use of a large language model (LLM) solely for language editing and grammatical refinement of the current manuscript. All scientific content, analysis, and interpretations presented herein are the sole responsibility of the authors.

### v3

@@
-CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning
+Denis Makhov1    Dmitriy Shopkhoev1,2 Magauiya Zhussip1    Ammar Ali1,2    Stamatios Lefkimmiatis1 Fundamental Research Center
 Abstract
-Post-training compression of large language models (LLMs) largely relies on low-rank weight approximation, which represents each column of a weight matrix in a shared low-dimensional subspace.
-While this is a computationally efficient strategy, the imposed structural constraint is rigid and can lead to a noticeable model accuracy drop.
-In this work, we propose CoSpaDi (Compression via Sparse Dictionary Learning), a novel training-free compression framework that replaces low-rank decomposition with a more flexible structured sparse factorization in which each weight matrix is represented with a dense dictionary and a column-sparse coefficient matrix.
-This formulation enables a union-of-subspaces representation: different columns of the original weight matrix are approximated in distinct subspaces spanned by adaptively selected dictionary atoms, offering greater expressiveness than a single invariant basis.
-Crucially, CoSpaDi leverages a small calibration dataset to optimize the factorization such that the output activations of compressed projection layers closely match those of the original ones, thereby minimizing functional reconstruction error rather than mere weight approximation.
-This data-aware strategy preserves better model fidelity without any fine-tuning under reasonable compression ratios.
-Moreover, the resulting structured sparsity allows efficient sparse–dense matrix multiplication and is compatible with post-training quantization for further memory and latency gains.
-We evaluate CoSpaDi across multiple Llama and Qwen models under per-layer and per-group settings at 205020-50% compression ratios, demonstrating consistent superiority over state-of-the-art data-aware low-rank methods both in accuracy and perplexity.
-Our results establish structured sparse dictionary learning as a powerful alternative to conventional low-rank approaches for efficient LLM deployment.
+Post-training compression of large language models (LLMs) often relies on low-rank weight approximations that represent each column of the weight matrix in a shared low-dimensional subspace.
+This strategy is computationally efficient but the underlying constraint can be overly rigid for heterogeneous projection weights and may incur avoidable accuracy loss.
+We propose CoSpaDi (Compression via Sparse Dictionary Learning), a training-free framework that replaces low-rank factorization with a structured sparse decomposition in which each weight matrix is represented as a dense dictionary multiplied by a column-sparse coefficient matrix.
+This yields a union-of-subspaces model: the columns of the weight matrix are represented as linear combinations of different subsets of dictionary atoms, improving expressiveness at a fixed parameter budget.
+CoSpaDi is calibration-guided: using a small calibration set, we optimize the factorization to minimize functional reconstruction error of layer outputs rather than weight-space error.
+An activation-derived Gram orthonormalization reformulates this data-aware objective into a standard dictionary learning problem on transformed weights, and we support both per-layer compression and cross-layer dictionary sharing within groups of similar projections.
+Across Llama and Qwen model families, CoSpaDi consistently improves the accuracy–compression and perplexity–compression trade-offs over state-of-the-art SVD-based baselines and strong structured pruning baselines at 20-40% compression ratios.
+The resulting structured sparsity enables sparse–dense computation and integrates with post-training quantization of the sparse coefficients.
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of applications, from conversational agents Brown et al.
-(2020); OpenAI (2023) to general-purpose reasoning systems Touvron et al.
-(2023a); Anil et al.
-(2023), owing to their ability to capture long-range dependencies through attention mechanisms Vaswani et al.
-(2017); Devlin et al.
-(2019).
-However, this success entails substantial memory and computational demands during both training and inference, which severely limits their deployment in resource-constrained environments.
-Various compression and acceleration techniques have been proposed over recent years to mitigate these challenges, including pruning Frankle & Carbin (2019); Sanh et al.
-(2020), quantization Dettmers et al.
-(2022); Xiao et al.
-(2023); Yao et al.
-(2022); Dettmers et al.
-(2023), and weight decomposition/factorization Denton et al.
-(2014); Hu et al.
-(2022); Liu et al.
-(2024); Zhang et al.
-(2023); Ma et al.
-(2019); Chen et al.
-(2018); Hsu et al.
-(2022); Yu & Wu (2023); Wang et al.
-(2025a); Han et al.
-(2015).
-Recently, there has been growing interest in effective compression of large models without retraining or fine-tuning, since this is often computationally prohibitive.
-A prominent line of work leverages low-rank approximations of weight matrices via Singular Value Decomposition (SVD).
-Early approaches such as DRONE Chen et al.
-(2021), FWSVD Hsu et al.
-(2022) and ASVD Yuan et al.
-(2023) introduced activation-aware truncation of singular values, while SVD-LLM Wang et al.
-(2025a) and its variant SVD-LLMv2 Wang et al.
-(2025b) proposed truncation-aware and optimized singular value selection strategies.
-In parallel, Basis Sharing Wang et al.
-(2024) explored cross-layer sharing of a common basis to further enhance compression efficiency by exploiting inter-layer redundancy.
-Despite their effectiveness, these methods are inherently constrained by the use of a single shared low-dimensional subspace, which may limit representational flexibility and prevent full exploitation of redundancies in LLM parameters.
-In this work, we argue that moving beyond the low-rank weight decomposition is a promising new direction.
-Specifically, we propose to adopt dictionary learning, a well-established paradigm in signal and image processing Aharon et al.
-(2006); Elad (2010), and apply it to LLM compression.
-Unlike low-rank approximations that confine all columns of a weight matrix to a shared linear subspace, dictionary learning enables a richer union-of-subspaces representation: each column is approximated using only a sparse subset of atoms from a learned dictionary, allowing different columns to reside in distinct subspaces spanned by different atom combinations.
-Such a representation better accommodates heterogeneous features and introduces additional flexibility to reduce the overall approximation error.
-To this end, we introduce CoSpaDi (Compression via Sparse Dictionary Learning), a training-free framework that applies dictionary learning to jointly estimate dictionaries and sparse coefficients used for compressing LLM weight matrices.
-Our contributions are as follows:
-  • We introduce dictionary learning with sparse coding as a new paradigm for LLM compression, addressing the limitation of using a single invariant basis for the weight approximation which is inherent in SVD-based methods.
-  • We demonstrate that CoSpaDi is effectively integrated with data-aware optimization, yielding state-of-the-art compression performance for a wide range of compression ratios, while being compatible with post-training quantization for further memory and latency gains.
-  • Through extensive experiments, we show that our approach is effective in both per-layer and cross-layer (shared dictionary) compression scenarios, consistently outperforming regular SVD, ASVD Yuan et al.
-(2023), SVD-LLM Wang et al.
-(2025a), and Basis Sharing Wang et al.
-(2024) strategies as well as state-of-the-art structure-pruning methods Ma et al.
-(2023); Shopkhoev et al.
-(2025).
+Large language models (LLMs) achieve strong performance across diverse tasks, from dialogue and instruction following (Brown et al., 2020; Achiam et al., 2023) to general-purpose reasoning (Touvron et al., 2023; Anil et al., 2023).
+Their effectiveness stems in part from their Transformer-based architectures that model long-range dependencies with attention (Vaswani et al., 2017; Devlin et al., 2019).
+At the same time, the scale that enables these capabilities makes LLMs expensive to store and run, creating a practical barrier to deployment on memory- and compute-constrained hardware.
+A broad literature addresses post-training compression and acceleration of LLMs, spanning pruning, quantization, knowledge distillation, and structured weight parameterizations such as matrix factorization (Frankle & Carbin, 2019; Dettmers et al., 2022; Hinton et al., 2015; Denton et al., 2014).
+Among training-free approaches, matrix factorization is particularly attractive because it yields explicit low-parameter surrogates for large projection matrices, and in practice the predominant choice is truncated SVD and its activation-aware variants that use a small calibration set to guide which components to keep and how to scale them (Chen et al., 2021).
+Cross-layer extensions further reduce overhead by sharing a common low-dimensional subspace across groups of layers (Wang et al., 2025a).
+Despite strong results, these methods ultimately approximate each matrix within a single shared low-dimensional subspace; for heterogeneous Transformer projections, this constraint can be unnecessarily restrictive and motivates richer factorizations.
+We study an alternative factorization family that relaxes this constraint.
+Instead of approximating a matrix with one global low-dimensional basis, we model it using a union of subspaces via sparse dictionary learning (Aharon et al., 2006; Elad, 2010): a learned dictionary of atoms is combined with column-sparse coefficients, allowing different columns to be reconstructed from different subsets of atoms.
+This representation is more flexible than a single shared subspace at the same parameter budget, and it is well aligned with the intuition that different output channels may depend on different latent features.
+Building on this idea, we propose CoSpaDi (Compression via Sparse Dictionary Learning), a training-free compression framework for Transformer projections.
+CoSpaDi learns dictionaries and sparse codes to approximate pretrained weight matrices, and is data-aware: from a small calibration set we construct an activation-derived Gram orthonormalization that reformulates functional output reconstruction into a standard dictionary learning problem on a transformed (activation-weighted) weight matrices.
+The resulting factorization yields structured sparsity that can be paired with post-training quantization of the sparse coefficients, and it naturally supports both per-layer compression and cross-layer dictionary sharing for groups of related projections.
+Contributions.
+  • Beyond low-rank for post-training compression.
+We introduce sparse dictionary learning as a compression paradigm for LLM weight matrices, replacing the single-subspace constraint of SVD-based factorization with a union-of-subspaces representation.
+  • Calibration-guided, training-free optimization.
+We integrate sparse dictionary learning with a data-aware objective via activation-derived Gram orthonormalization, yielding a tractable transformed problem that can be optimized efficiently with alternating sparse coding and dictionary updates, without gradient-based fine-tuning.
+  • Empirical gains across settings.
+Across multiple Llama and Qwen models and a range of compression ratios, CoSpaDi improves the quality–compression trade-off over strong activation-aware SVD baselines (including SVD-LLM (Wang et al., 2025c) in the per-layer setting and Basis Sharing (Wang et al., 2025a) in the grouped setting), and is competitive with recent structured pruning methods (Ma et al., 2023; Shopkhoev et al., 2025).
 2 Related Work
-Research on Transformer compression covers pruning, quantization, distillation, and matrix factorization. Below we provide a brief overview of the recent progress in these directions.
-Early work on pruning showed substantial redundancy in deep nets Han et al.
-(2015); Frankle & Carbin (2019); Sanh et al.
-(2020).
-For LLMs, SparseGPT performs one-shot post-training pruning and achieves high sparsity with small quality drop Frantar & Alistarh (2023).
-Wanda reduces overhead further with simple activation-based rules Sun et al.
-(2023).
-LLM-Pruner removes blocks using gradient-guided criteria and recovers accuracy with brief adapter tuning Ma et al.
-(2023), while ReplaceMe substitutes multiple transformer blocks with a single linear transformation Shopkhoev et al.
-(2025).
-Quantization reduces precision to reduce memory consumption and accelerate inference.
-For weight-only post-training quantization, LLM.int8 enables INT8 inference via vector-wise quantization with a path for outliers Dettmers et al.
-(2022); GPTQ/OPTQ minimize output error using Hessian to reach 3–4 bit weights Frantar et al.
-(2022; 2023); AWQ derives scales from activations and protects a small set of salient weights Lin et al.
-(2024), and SpQR stores outliers at higher precision to maintain quality at 3–4 bits Dettmers et al.
-(2024).
-For quantizing both weights and activations, SmoothQuant rebalances channel ranges to achieve accurate W8A8 across matrix multiplications Xiao et al.
-(2023), while QuaRot applies orthogonal rotations to suppress outliers, enabling end-to-end 4-bit inference including the KV cache Ashkboos et al.
-(2024).
-The general framework on knowledge distillation by Hinton et al.
-and sequence-level distillation by Kim and Rush established strong baselines Hinton et al.
-(2015); Kim & Rush (2016).
-DistilBERT demonstrates task-agnostic compression for Transformers Sanh et al.
-(2019); TinyBERT uses a two-stage distillation procedure Jiao et al.
-(2020); MobileBERT introduces a teacher-guided compact architecture Sun et al.
-(2020); Patient-KD leverages multi-layer hints Sun et al.
-(2019); MiniLM and MiniLMv2 distill self-attention relations Wang et al.
-(2020; 2021).
-BERT-of-Theseus compresses models via progressive module replacement Xu et al.
-(2020).
-Orca and Orca 2 show that richer teacher signals such as explanations can improve reasoning in smaller students Mukherjee et al.
-(2023); Mitra et al.
-(2023).
-Classic low-rank approximations reduce parameters and FLOPs with limited loss Denton et al.
-(2014).
-DRONE considers an objective related to the output activation for the weight approximation, Chen et al.
-(2021).
-FWSVD introduces Fisher-weighted reconstruction to emphasize important directions Hsu et al.
-(2022), while ASVD adapts truncation using activation-aware transforms and layer sensitivity Yuan et al.
-(2023).
-SVD-LLM utilizes truncation-aware whitening , while SVD-LLM v2 optimizes budget allocation across layers Wang et al.
-(2025a; b).
-Basis Sharing reuses one part of low-rank factorization across layers and learns layer-specific coefficients to exploit inter-layer redundancy Wang et al.
-(2024).
-Dictionary learning and sparse coding.
-Dictionary learning factorizes original weight into dictionary and sparse codes, yielding a union-of-subspaces model with strong results in vision and image compression Engan et al.
-(1999); Aharon et al.
-(2006); Mairal et al.
-(2010); Gregor & LeCun (2010); Elad (2010).
-In NLP, GroupReduce explores block-wise low-rank/dictionary-style compression Chen et al.
-(2018), and tensorized Transformers factorize attention and embeddings Ma et al.
-(2019).
-Recent work also targets the KV-cache by learning universal dictionaries and decoding with OMP during inference Kim et al.
-(2024).
-Complementary to these directions, a recent work proposes cross-layer weight sharing via a matrix-based dictionary learning formulation for Transformer attention, directly exploiting inter-layer redundancy Zhussip et al.
-(2025).
-Cross-layer/shared-basis schemes in attention are thus closely related Wang et al.
-(2024).
-A comprehensive, training-free approach to weight-space dictionary learning for LLMs — at both per-layer and cross-layer levels — remains underexplored and is the focus of this work.
-Figure 1: Left side: weight factorization methods using low-rank decomposition.
-Low-rank approximation decomposes a matrix into two dense matrices of lower rank.
-Right side: proposed CoSpaDi.
-A dictionary of $k$ atoms and a column-sparse coefficient matrix are employed.
-No restrictions on size of $k$ (undercomplete : $k<d_{1}$, complete: $k=d_{1}$ or overcomplete : $k>d_{1}$ dictionaries are possible), while sparsity is defined by $s$ non-zero elements per column of the coefficient matrix.
-3 Proposed Method
-In this section, we first provide a conceptual link between low-rank weight approximation and dictionary learning.
-Then, we introduce the proposed Sparse Dictionary Learning (SDL) strategy for LLM compression covering all practical aspects including data-aware SDL, dictionary-sharing among layers, compression ratios and inference complexity.
-3.1 Low-Rank Weight Approximation
-A widely adopted strategy to reduce the parameter count in LLMs is to approximate weight matrices $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ — representing the parameters of network layers — with matrices of reduced rank $r<\min(d_{1},d_{2})$.
-Such a low-rank approximation can be derived as follows:
-$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\operatorname*{rank}\left(\tilde{\mathbf{W}}\right)=r}\left\|\mathbf{W}-\tilde{\mathbf{W}}\right\|_{F}.$ (1)
-According to the Eckart–Young–Mirsky theorem Eckart & Young (1936), the orthogonal projection to the space of $r$-rank matrices admits an analytical solution.
+Post-training compression for Transformers is commonly organized around what is preserved (weights, activations, or end-to-end behavior) and what structure is imposed (sparsity, low precision, low rank, or shared components).
+We review the most relevant directions through this lens, and position CoSpaDi as a functional, calibration-guided method that imposes a union-of-subspaces structure via sparse dictionary learning.
+Post-training pruning and practical sparsity.
+A central theme in recent LLM compression is to optimize a compressed representation to match layer outputs on a small calibration set, rather than minimizing weight-space error.
+This perspective appears prominently in one-shot unstructured pruning methods such as SparseGPT (Frantar & Alistarh, 2023) and its simplified activation-driven variants (e.g., Wanda) (Sun et al., 2024), which induce high sparsity with limited degradation by explicitly targeting output preservation.
+However, unstructured sparsity does not automatically translate into practical compression or speedups: sparse storage formats typically require additional metadata (e.g., indices in CSR/CSC-like representations), which can substantially reduce effective memory gains at moderate sparsity, and irregular access patterns often make sparse matmuls memory-bound and kernel-limited (Han et al., 2015; Wang, 2020).
+In practice, inference acceleration usually benefits most from structured sparsity patterns and dedicated kernel/hardware support (e.g., 2:4 sparse tensor cores) (Mishra et al., 2021; NVIDIA, 2020).
+Complementary structured approaches remove or replace higher-level components (blocks/layers) using importance criteria and brief recovery steps (Ma et al., 2023; Shopkhoev et al., 2025).
+Our work follows the same calibration-guided spirit, but instead of removing weights, we seek a structured factorization that can represent heterogeneous columns more flexibly than a shared low-dimensional subspace.
+Quantization and equivalent transformations.
+Quantization reduces memory footprint and can improve throughput, but accuracy at low bit-width often hinges on handling outliers and reshaping activation/weight distributions.
+Weight-only PTQ methods include vector-wise quantization with outlier routing (Dettmers et al., 2022), Hessian-aware schemes such as GPTQ/OPTQ (Frantar et al., 2023a; b), and activation-aware scaling/outlier protection (Lin et al., 2024; Dettmers et al., 2024).
+For weight-and-activation quantization, methods such as SmoothQuant (Xiao et al., 2023) and rotation-based transforms (e.g., QuaRot) (Ashkboos et al., 2024) improve quantizability by rebalancing channel ranges or suppressing outliers.
+These techniques are complementary to structural compression: in our setting, the sparse coefficients produced by CoSpaDi can be quantized post hoc to further reduce memory and potentially improve practical efficiency.
+Low-rank factorization and shared subspaces.
+A widely used post-training alternative to pruning is matrix factorization, where each projection is approximated with a low-rank surrogate (Denton et al., 2014).
+Activation-aware and functional variants include DRONE (Chen et al., 2021), Fisher-weighted reconstruction (FWSVD) (Hsu et al., 2022), and adaptive truncation with activation transforms (ASVD) (Yuan et al., 2023).
+More recent SVD-LLM methods incorporate truncation-aware whitening and improved allocation strategies across layers (Wang et al., 2025c; b; Qinsi et al., 2025).
+In parallel, Basis Sharing (Wang et al., 2025a) ties factors across layers to exploit inter-layer redundancy.
+While effective, these approaches still represent each matrix in a single shared subspace (possibly partially shared across layers).
+Our method targets the same post-training regime, but replaces the single-subspace constraint with a union-of-subspaces representation enabled by sparse coding.
+Dictionary learning and sparse representations.
+Dictionary learning and sparse coding have a long history in signal processing and vision (Engan et al., 1999; Aharon et al., 2006; Mairal et al., 2009; Gregor & LeCun, 2010; Elad, 2010), where they provide expressive, parsimonious representations that adaptively select a small subset of atoms per example.
+In model compression, related ideas appear in structured decompositions such as GroupReduce (Chen et al., 2018) and tensorized factorization schemes (Ma et al., 2019).
+Recent work has also applied learned dictionaries to other Transformer components, e.g., KV-cache compression via sparse decoding (Kim et al., 2025), and cross-layer sharing formulations specialized to attention (Zhussip et al., 2025), which connects to broader shared-basis approaches (Wang et al., 2025a).
+In contrast, our focus is a training-free, calibration-guided dictionary learning framework for weight compression that (i) explicitly optimizes a functional objective through activation-aware Gram orthonormalization, and (ii) supports both per-layer compression and cross-layer dictionary sharing within a unified procedure.
+Figure 1: Left: low-rank factorization represents $\mathbf{W}$ in a single $r$-dimensional subspace via two dense factors.
+Right: CoSpaDi represents $\mathbf{W}$ as $\mathbf{D}\mathbf{S}$ where $\mathbf{D}$ is a dictionary of $k$ atoms and $\mathbf{S}$ is column-sparse (at most $s$ nonzeros per column), yielding a union-of-subspaces model.
+Dictionaries may be undercomplete ($k<d_{1}$), complete ($k=d_{1}$), or overcomplete ($k>d_{1}$).
+3 Method
+We first formulate post-training compression objectives in weight space and activation space.
+We then reinterpret truncated SVD through the lens of PCA as a “basis times coefficients” model, and introduce sparse dictionary learning as a more expressive union-of-subspaces alternative.
+Finally, we describe CoSpaDi and its grouped extension, together with compression and complexity estimates.
+3.1 Problem formulation: weight-space vs. activation-space objectives
+Consider a pretrained linear projection with weight matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ and input $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ (a batch of $N$ calibration activations).
+The layer output is $\mathbf{Z}=\mathbf{X}\mathbf{W}\in\mathbb{R}^{N\times d_{2}}$.
+Post-training compression replaces $\mathbf{W}$ by a structured approximation $\tilde{\mathbf{W}}$ that reduces storage and ideally improves inference efficiency.
+A common starting point is weight-space reconstruction:
+$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}\in\mathcal{C}}\|\mathbf{W}-\tilde{\mathbf{W}}\|_{F}^{2},$ (1)
+where $\mathcal{C}$ encodes a chosen structure (e.g., rank-$r$, sparsity, quantization, or a parametric factorization).
+For LLM projections, however, preserving weights is often a weak proxy for preserving model behavior. A more directly relevant goal is activation-space (functional) reconstruction on calibration data:
+$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}\in\mathcal{C}}\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\|_{F}^{2}.$ (2)
+Objective (2) is used (implicitly or explicitly) across many calibration-guided post-training methods, and motivates the activation-aware design of CoSpaDi.
+3.2 Truncated SVD as PCA: basis $\times$ coefficients
+Low-rank approximation chooses $\mathcal{C}=\{\tilde{\mathbf{W}}:\mathrm{rank}(\tilde{\mathbf{W}})\leq r\}$ and solves (1):
+$\tilde{\mathbf{W}}^{\mathrm{LR}}=\operatorname*{arg\,min}_{\mathrm{rank}(\tilde{\mathbf{W}})\leq r}\|\mathbf{W}-\tilde{\mathbf{W}}\|_{F}^{2}.$ (3)
+According to the Eckart–Young–Mirsky theorem (Eckart & Young, 1936), the orthogonal projection to the space of $r$-rank matrices admits an analytical solution.
 Specifically, if $\mathbf{W}$ admits the singular value decomposition $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, with $\mathbf{U}\in\mathbb{R}^{d_{1}\times k}$, $\bm{\Sigma}\in\mathbb{R}^{k\times k}$ and $\mathbf{V}\in\mathbb{R}^{d_{2}\times k}$ with $k=\min\left(d_{1},d_{2}\right)$, then the minimizer of Eq.
-(1) can be expressed as: $\tilde{\mathbf{W}}^{\star}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$, where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ contains the first $r$ left singular vectors, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ holds the top-$r$ singular values, and $\mathbf{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ contains the first $r$ right singular vectors of $\mathbf{W}$, respectively.
-There is a deep connection between this problem and the principal component analysis (PCA) Bishop & Nasrabadi (2006).
+(3) can be expressed as: $\tilde{\mathbf{W}}^{\mathrm{LR}}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$, where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ contains the first $r$ left singular vectors, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ holds the top-$r$ singular values, and $\mathbf{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ contains the first $r$ right singular vectors of $\mathbf{W}$, respectively.
+There is a deep connection between this problem and the principal component analysis (PCA) (Bishop & Nasrabadi, 2006).
 Specifically, consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$.
 We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$.
-The optimal basis and coefficients can be found by minimizing the total approximation error:
-$\mathcal{J}=\sum\limits_{j=1}^{d_{2}}\left\|\mathbf{w}_{j}-\sum\limits_{i=1}^{r}c_{i,j}\bm{b}_{i}\right\|_{2}^{2}\equiv\left\|\mathbf{W}-\mathbf{B}\mathbf{C}\right\|_{F}^{2},$ (2)
-subject to the orthogonality constraint $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}$, where $\mathbf{B}=\begin{bmatrix}\bm{b}_{1},\ldots,\bm{b}_{r}\end{bmatrix}\in\mathbb{R}^{d_{1}\times r}$ is the basis matrix, $\mathbf{C}\in\mathbb{R}^{r\times d_{2}}$ is the coefficient matrix with entries $c_{i,j}$, and $\mathbf{I}\in\mathbb{R}^{r\times r}$ is the identity matrix.
-The optimal pair $\left(\mathbf{B}^{*},\mathbf{C}^{*}\right)$ is obtained by solving:
-$\mathbf{B}^{*},\mathbf{C}^{*}=\operatorname*{arg\,min}_{\mathbf{B},\mathbf{C}}\mathcal{J}\,\,\mbox{s.t}\,\,\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}.$ (3)
-The solution is given by $\mathbf{B}^{*}=\mathbf{U}_{r}$ and $\mathbf{C}^{*}=\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$ (we refer to Appendix A.1 for the proof).
-This result highlights that in the low-rank factorization $\mathbf{W}\approx\mathbf{B}\mathbf{C}$, the matrix $\mathbf{B}$ corresponds to the basis of the shared $r$-dimensional subspace in which the columns of $\mathbf{W}$ approximately lie, and $\mathbf{C}$ contains their coordinate vectors (coefficients) with respect to that basis.
-3.2 Sparse Dictionary Learning
-Motivated by this interpretation, we propose an alternative strategy for compressing the weights $\mathbf{W}$, based on the sparse dictionary learning methodology.
-Specifically, rather than modeling each column of the weight matrix as a linear combination of basis vectors $\bm{b}_{i}$, we aim to learn a dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and approximate each column $\mathbf{w}_{j}$ as a linear combination of dictionary atoms $\mathbf{d}_{i}\in\mathbb{R}^{d_{1}}\text{ with }i=1,\ldots,k$.
-Although the two strategies appear similar, they differ in two key respects.
-First, in dictionary learning, unlike low-rank approximation, only a subset of atoms is used to represent each weight vector $\mathbf{w}_{j}$.
-This principle, known in the literature as sparse coding Lee et al.
-(2006), is motivated by the observation that learned dictionaries typically yield sparse representations: each signal (column) activates only a few atoms.
-In other words, instead of approximating the weight matrix $\mathbf{W}$ as the product of two dense matrices, the orthogonal basis $\mathbf{B}$ and the coefficient matrix $\mathbf{C}$, we approximate it as the product of a dense dictionary matrix $\mathbf{D}$ and a column-wise sparse coefficient matrix $\mathbf{S}$, where each column $\mathbf{s}_{j}$ indicates which atoms contribute to the representation of $\mathbf{w}_{j}$, and with what weights.
-Second, and of significant practical consequence, we do not impose orthogonality constraints on the dictionary atoms.
-This grants greater representational flexibility, allowing the model to adapt more effectively to the intrinsic structure of the weight matrix.
-Finally, under the proposed strategy, rather than enforcing a shared low-dimensional subspace for all columns, each weight vector $\mathbf{w}_{j}$ is represented within its own low-dimensional subspace spanned by the atoms it activates.
-Consequently, the entire matrix $\mathbf{W}$ is modeled as a union of low-dimensional subspaces, introducing additional representational capacity and enabling lower approximation error compared to a single shared subspace.
-Formally, the dictionary learning problem can be expressed as:
-$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{W}-\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2},$ (4)
-where $\mathbf{s}_{j}$ denotes the $j$-th column of the sparse coefficient matrix $\mathbf{S}$, and $\left\|\cdot\right\|_{0}$ denotes the $\ell_{0}$ pseudo-norm, which counts the number of nonzero entries in a vector.
-The optimization problem in Eq.
-(4) is NP-hard in its original form and admits no closed-form solution.
-Nevertheless, it has been extensively studied, and several efficient algorithms have been proposed to obtain high-quality approximate solutions.
-Among these, K-SVD Aharon et al.
-(2006) is a well-established algorithm widely adopted across diverse applications.
-K-SVD alternates between two main steps: (a) Sparse coding: Each signal $\mathbf{w}_{j}$ is approximated as a sparse linear combination of dictionary atoms (typically via orthogonal matching pursuit or LASSO).
-(b) Dictionary update: Each atom $\mathbf{d}_{i}$ is refined to better fit the data while preserving the sparsity pattern in $\mathbf{S}$.
-This is performed sequentially per atom: the algorithm identifies the subset of signals that activate $\mathbf{d}_{i}$, computes the residual error over those signals, and applies a rank-1 SVD to jointly update the atom and its associated coefficients.
-3.3 CoSpadi: Activation-Aware SDL
-In several recent works, including Chen et al.
-(2021), it has been observed that LLM weight matrices generally do not exhibit intrinsic low-rank structure.
-Consequently, direct low-rank approximation of the weights often leads to significant performance degradation.
-However, under the hypothesis that the input latent features to a given layer reside in a low-dimensional subspace, it remains feasible and often effective to approximate the corresponding layer weights as low-rank.
-This approach is well-motivated: the goal is not to preserve the weights in isolation, but rather to maintain the fidelity of the activations they produce when applied to low-dimensional inputs.
-Within this framework, an effective approximation of $\mathbf{W}$ can be obtained by minimizing the output error:
-$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\right\|_{F}\quad\text{s.t.}\quad\operatorname*{rank}(\mathbf{X}\tilde{\mathbf{W}})=r,$ (5)
-where $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ is a matrix of $N$ calibration input vectors $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d_{1}}\right\}_{i=1}^{N}$.
-This minimization admits the analytical solution: $\tilde{\mathbf{W}}=\mathbf{L}^{-1}\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}=\mathbf{B}\mathbf{C}$, where $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is a non-singular transformation matrix derived from the calibration data (see Appendix A.2 for the derivation).
-Motivated by this data-aware perspective, we propose adapting our sparse dictionary learning framework accordingly.
-Rather than directly approximating $\mathbf{W}$ as $\tilde{\mathbf{W}}=\mathbf{D}\mathbf{S}$, we instead minimize the reconstruction error of the output activations $\mathbf{Z}=\mathbf{X}\mathbf{W}$:
-$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (6)
-To simplify this optimization, let $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}\in\mathbb{R}^{N\times d_{1}}$ be a column-orthogonal matrix, i.e., $\mathbf{Y}^{\mathsf{T}}\mathbf{Y}=\mathbf{I}_{d_{1}}$, obtained by applying a linear transformation $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ to $\mathbf{X}$.
-We assume $N\geq d_{1}$ and that $\mathbf{X}$ has full column rank — conditions typically satisfied with sufficient calibration data.
-The matrix $\mathbf{L}$ can be computed via QR decomposition, SVD of $\mathbf{X}$, or Cholesky/eigen-decomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
-Introducing the auxiliary variables $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}$ and $\mathbf{D}_{L}=\mathbf{L}\mathbf{D}$, and noting that the Frobenius norm is invariant under left-multiplication by a column-orthogonal matrix, we can reformulate Eq.
-(6) as:
-$\mathbf{D}_{L}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D}_{L},\mathbf{S}}\left\|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (7)
-Since $\mathbf{L}$ is invertible, the final structured approximation of the original weights is given by:
-$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S},\quad\text{where}\quad\mathbf{D}_{a}=\mathbf{L}^{-1}\mathbf{D}_{L}$ (8)
-is the activation-aware learned dictionary. Pseudocode for the full procedure is provided in Appendix A.3.
-Cross-Layer SDL
-Although most compression techniques focus on intra-layer optimizations, the repetitive layered structure of LLMs suggests the presence of significant inter-layer redundancy.
-To exploit this, we propose sharing a single dictionary across weight matrices of the same type from multiple layers.
-Specifically, let $\mathcal{G}$ denote a group of layer indices (e.g., all attention projection layers or all FFN layers).
-We form a grouped weight matrix by horizontally concatenating the corresponding layer weights: $\mathbf{W}_{G}=\begin{bmatrix}\mathbf{W}_{\ell_{1}},&\mathbf{W}_{\ell_{2}},&\ldots,&\mathbf{W}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{d_{1}\times(d_{2}\cdot L)},$ where each $\mathbf{W}_{\ell}\in\mathbb{R}^{d_{1}\times d_{2}}$, and $L=|\mathcal{G}|$ is the number of layers in the group.
-We then apply the same activation-aware dictionary learning procedure as before, but now to $\mathbf{W}_{G}$.
-To ensure the approximation remains data-aware across all layers in the group, we construct a grouped calibration matrix by vertically stacking the corresponding input batches: $\mathbf{X}_{G}=\begin{bmatrix}\mathbf{X}_{\ell_{1}};&\mathbf{X}_{\ell_{2}};&\ldots;&\mathbf{X}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{(N\cdot L)\times d_{1}}$ where each $\mathbf{X}_{\ell}\in\mathbb{R}^{N\times d_{1}}$ is the calibration input for layer $\ell$.
-The transformation matrix $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is then computed from $\mathbf{X}_{G}$ (e.g., via QR or Cholesky decomposition), ensuring that $\mathbf{Y}_{G}=\mathbf{X}_{G}\mathbf{L}^{-1}$ is column-orthogonal.
-The optimization proceeds by solving Eq.
-(7) with $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}_{G}$, yielding a shared dictionary $\mathbf{D}_{a}$ and a sparse coefficient matrix $\mathbf{S}_{G}\in\mathbb{R}^{k\times(d_{2}\cdot L)}$.
-This approach, inspired by Wang et al.
-(2024), reduces memory overhead by amortizing the dictionary cost across multiple layers, while preserving activation fidelity through data-aware calibration.
-Each layer’s compressed weights are then recovered by slicing the corresponding block from $\tilde{\mathbf{W}}_{G}=\mathbf{D}_{a}\mathbf{S}_{G}$.
-Compression ratio.
-For the low-rank setting, the compression ratio is $\gamma^{\mathrm{LR}}\coloneqq 1-\tfrac{r(d_{1}+d_{2})}{d_{1}d_{2}}$, where $r$ is the retained rank.
-For CoSpaDi, we store a dictionary and sparse codes plus a binary mask (details are provided in section A.4), giving
-$\gamma^{\mathrm{SD}}\coloneqq 1-\tfrac{\;\overbrace{d_{1}k}^{\text{dict. (bf16)}}+\overbrace{sd_{2}}^{\text{codes (bf16)}}+\overbrace{(kd_{2})/16}^{\text{mask (1 bit/entry)}}\;}{d_{1}d_{2}}.$ (9)
-Unlike the low-rank case, $\gamma^{\mathrm{SD}}$ depends on two knobs — the number of atoms $k$ and sparsity $s$ — allowing us to trade off model capacity and storage at fixed budget.
-We parameterize this trade-off by the ratio $\rho\coloneqq k/s$, which uniquely determines $(k,s)$ at a target $\gamma^{\mathrm{SD}}$:
+The optimal basis $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$ and coefficients $\mathbf{C}\in\mathbb{R}^{r\times d_{2}}$ can be found by minimizing the total approximation error:
+$\min_{\mathbf{B},\mathbf{C}}\ \|\mathbf{W}-\mathbf{B}\mathbf{C}\|_{F}^{2}\quad\text{s.t.}\ \mathbf{B}^{\top}\mathbf{B}=\mathbf{I}.$ (4)
+The optimum is $\mathbf{B}^{\star}=\mathbf{U}_{r}$ and $\mathbf{C}^{\star}=\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$ (see Appendix A.1 for the proof).
+Therefore, truncated SVD realizes a single shared $r$-dimensional subspace: every column $\mathbf{w}_{j}$ is expressed in the same basis $\mathbf{B}$ via dense coordinates (the $j$-th column of $\mathbf{C}$).
+Although Eq.
+(3) minimizes reconstruction error in weight space, several works (e.g., (Chen et al., 2021)) note that LLM projection matrices often lack strong intrinsic low-rank structure, so directly approximating $\mathbf{W}$ can lead to noticeable degradation.
+A common alternative is to optimize functional reconstruction on calibration activations, motivated by the hypothesis that layer inputs lie in a lower-dimensional subspace and that preserving the induced outputs is more relevant than preserving weights in isolation.
+Concretely, this leads to minimizing $\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\|_{F}^{2}$ for calibration inputs $\mathbf{X}$ (Chen et al., 2021; Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2025c).
+In this setting, the optimal low-rank approximation can still be obtained in closed form after applying an activation-derived Gram orthonormalization; we provide the full derivation and solution in Appendix A.2.
+3.3 Sparse dictionary learning: union of subspaces
+Dictionary learning keeps the “basis times coefficients” form but changes the modeling assumptions.
+We replace the orthonormal basis $\mathbf{B}$ by a (generally non-orthogonal) dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and enforce column sparsity in the coefficient matrix $\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$:
+$\min_{\mathbf{D},\mathbf{S}}\ \|\mathbf{W}-\mathbf{D}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j,$ (5)
+where $\mathbf{s}_{j}$ is the $j$-th column of $\mathbf{S}$, $\left\|\cdot\right\|_{0}$ denotes the $\ell_{0}$ pseudo-norm, and $s$ indicates the sparsity level.
+Compared to low-rank factorization, Eq.
+(5) differs in two crucial ways: (i) coefficients are sparse, so each column activates only $s$ atoms, and (ii) no orthogonality constraint is imposed on $\mathbf{D}$, allowing more flexible representations.
+As a result, columns can live in different subspaces spanned by different subsets of atoms: $\mathbf{W}$ is represented using a union of subspaces model rather than a single shared subspace.
+3.4 CoSpaDi: activation-aware sparse dictionary learning for LLM projections
+Motivated by the observation that activations, not projection weights, exhibit low-rank structure, we apply dictionary learning to LLM projection matrices under the activation-space objective (2).
+Concretely, we choose the constraint set $\mathcal{C}=\{\mathbf{D}\mathbf{S}:\mathbf{D}\in\mathbb{R}^{d_{1}\times k},\ \mathbf{S}\in\mathbb{R}^{k\times d_{2}},\ \|\mathbf{s}_{j}\|_{0}\leq s\}$ and solve
+$\min_{\mathbf{D},\mathbf{S}}\ \|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{D}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j.$ (6)
+Gram orthonormalization transform.
+To simplify Eq.
+(6), consider the Gram matrix $\mathbf{G}\coloneqq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{d_{1}\times d_{1}}$.
+We compute a Gram orthonormalization matrix $\mathbf{L}$ from the (uncentered) second moment $\mathbf{G}$ such that
+$\mathbf{Y}\coloneqq\mathbf{X}\mathbf{L}^{-1}\quad\text{satisfies}\quad\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{I},$ (7)
+i.e., $\mathbf{Y}$ has orthogonal columns.
+In practice, $\mathbf{L}$ can be obtained via QR factorization of $\mathbf{X}$, or from a factorization of $\mathbf{G}$ (e.g., Cholesky when $\mathbf{G}$ is symmetric positive definite, and an SVD/eigendecomposition-based otherwise).
+Next, we define the transformed variables $\mathbf{W}_{L}\coloneqq\mathbf{L}\mathbf{W}$ and $\mathbf{D}_{L}\coloneqq\mathbf{L}\mathbf{D}$.
+Using the invariance of the Frobenius norm under left multiplication by a column-orthogonal matrix, Eq.
+(6) reduces to
+$\min_{\mathbf{D}_{L},\mathbf{S}}\ \|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j.$ (8)
+After solving (8), we map back to the original parameterization:
+$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S},\qquad\mathbf{D}_{a}=\mathbf{L}^{-1}\mathbf{D}_{L}.$ (9)
+Thus, calibration data enters only through the “whitening” transform $\mathbf{L}$, while the core optimization reduces to standard sparse dictionary learning on $\mathbf{W}_{L}$.
+Optimization via alternating minimization.
+We approximately solve (8) by alternating between sparse coding and dictionary update:
+  • Sparse coding.
+For each column $\mathbf{w}_{L,j}$ of $\mathbf{W}_{L}$, we compute a sparse code $\mathbf{s}_{j}$ with at most $s$ nonzeros by approximately solving $\min_{\mathbf{s}_{j}}\|\mathbf{w}_{L,j}-\mathbf{D}_{L}\mathbf{s}_{j}\|_{2}^{2}\ \text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s$, using orthogonal matching pursuit (OMP) (Tropp & Gilbert, 2007).
+  • Dictionary update.
+With sparse supports fixed, we update the dictionary.
+Two common options are: (i) MOD (method of optimal directions), which updates all atoms jointly via a global least-squares step (Engan et al., 1999), and (ii) K-SVD, which updates atoms sequentially by solving a rank-1 approximation on a restricted residual (Aharon et al., 2006).
+Concretely, for atom $i$ let $\Omega_{i}=\{j:\ (\mathbf{s}_{j})_{i}\neq 0\}$ be the set of columns using atom $i$ and form the restricted residual $\mathbf{E}_{i}=\big(\mathbf{W}_{L}-\sum_{p\neq i}\mathbf{d}_{p}\mathbf{s}_{p,:}\big)_{[:,\Omega_{i}]}$.
+K-SVD updates $\mathbf{d}_{i}$ and $\mathbf{s}_{i,\Omega_{i}}$ via the top rank-1 component of $\mathbf{E}_{i}$, which can be computed exactly by SVD or approximated with truncated/randomized low-rank routines (Halko et al., 2011) or power iterations (Golub & Van Loan, 2013).
+In practice, both steps admit efficient implementations, e.g., Batch-OMP for sparse coding and approximate rank-1 updates for K-SVD, substantially reducing runtime and memory overhead (Rubinstein et al., 2008).
+Across our ablations, K-SVD with power-iteration rank-1 updates as well as Batch-OMP provides the strongest performance under a reasonable wall-clock budget.
+The full procedure is summarized in Appendix 1.
+3.5 Grouped / shared-dictionary extension
+To exploit inter-layer redundancy, we optionally share a dictionary across a group of projections of the same type.
+Let $\mathcal{G}=\{\ell_{1},\ldots,\ell_{L}\}$ be a set of layers with weights $\mathbf{W}_{\ell}\in\mathbb{R}^{d_{1}\times d_{2}}$.
+We form a grouped matrix by concatenation
+$\mathbf{W}_{\mathcal{G}}=\big[\mathbf{W}_{\ell_{1}},\ldots,\mathbf{W}_{\ell_{L}}\big]\in\mathbb{R}^{d_{1}\times(Ld_{2})}.$ (10)
+To preserve functional fidelity across the group, we stack calibration inputs $\mathbf{X}_{\mathcal{G}}=\big[\mathbf{X}_{\ell_{1}}^{\top},\ldots,\mathbf{X}_{\ell_{L}}^{\top}\big]^{\top}\in\mathbb{R}^{(LN)\times d_{1}}$, compute a Gram orthonormalization matrix $\mathbf{L}$ from $\mathbf{X}_{\mathcal{G}}$, and solve (8) on $\mathbf{W}_{\mathcal{G},L}=\mathbf{L}\mathbf{W}_{\mathcal{G}}$ to obtain a shared dictionary and block-structured sparse codes.
+Each layer is then recovered by slicing the corresponding coefficient block.
+This approach, inspired by Basis Sharing (Wang et al., 2025a), reduces memory overhead by amortizing the dictionary cost across multiple layers, while preserving activation fidelity through data-aware calibration.
+Each layer’s compressed weights are then recovered by slicing the corresponding coefficient block.
+Specifically, if $\mathbf{S}_{\mathcal{G}}\in\mathbb{R}^{k\times(Ld_{2})}$ are the learned codes, then the codes for layer $\ell_{t}$ are $\mathbf{S}_{\ell_{t}}=\mathbf{S}_{\mathcal{G}}[:,(t-1)d_{2}:td_{2}]$.
+3.6 Compression ratio
+For low-rank factorization with rank $r$, storing two dense factors costs $r(d_{1}+d_{2})$ parameters, leading to a compression ratio:
+$\gamma^{\mathrm{LR}}\coloneqq 1-\frac{r(d_{1}+d_{2})}{d_{1}d_{2}}.$ (11)
+For CoSpaDi, we store a dense dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and sparse coefficients $\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$.
+With $s$ nonzeros per column, the number of stored coefficient values is $sd_{2}$ (plus indices/mask; see Appendix A.4 for exact accounting).
+Thus, we can write
+$\gamma^{\mathrm{SD}}\coloneqq 1-\frac{\,\overbrace{d_{1}k}^{\text{dict.}}+\overbrace{sd_{2}}^{\text{values}}+\overbrace{(kd_{2})/16}^{\text{mask}}\,}{d_{1}d_{2}}.$ (12)
+Unlike low-rank compression, $\gamma^{\mathrm{SD}}$ depends on two knobs $(k,s)$, enabling additional flexibility at a fixed storage budget.
+We parameterize the trade-off via $\rho\coloneqq k/s$, which uniquely determines $(k,s)$ at a target $\gamma^{\mathrm{SD}}$:
 $k=\frac{(1-\gamma^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}+\tfrac{d_{2}}{16}\,},\qquad s=\tfrac{k}{\rho}.$ (13)
-Inference Complexity
-In terms of computational efficiency, low-rank and dictionary learning exhibit distinct inference-time complexity profiles: the former has a cost of $Tr(d_{1}+d_{2})$, whereas the latter – when exploiting sparsity and reusing inner products – achieves $Td_{1}K_{\text{active}}+Tsd_{2}$ (see section A.5), potentially yielding superior efficiency under favorable sparsity patterns.
-Although both methods share identical theoretical complexity under matched compression ratios, practical inference latency varies significantly due to factors such as the number of active atoms, indexing overhead, and hardware-specific kernel efficiency.
+3.7 Inference complexity
+In terms of computational efficiency, low-rank and dictionary learning exhibit distinct inference-time complexity profiles: the former has a cost of $Nr(d_{1}+d_{2})$, whereas the latter – when exploiting sparsity and reusing inner products – achieves $Nd_{1}K_{\text{active}}+Nsd_{2}$, potentially yielding superior efficiency under favorable sparsity patterns.
+Although both methods share identical theoretical complexity under matched compression ratios (Appendix A.5), practical inference latency varies significantly due to factors such as the number of active atoms, indexing overhead, and hardware-specific kernel efficiency.
 Further details regarding complexity derivations are provided in Appendix A.5.
 4 Experiments
-In this section, we present our experimental setup.
-We first compare CoSpaDi with low-rank baselines in the per-layer setting, where each linear weight matrix is compressed independently.
-We then investigate cross-layer sharing, analyze the impact of coefficient matrix quantization, and conduct ablation studies on the $ks$-ratio.
+We evaluate CoSpaDi in two regimes: per-layer compression, where each projection matrix is compressed independently, and grouped (cross-layer) compression, where a dictionary is shared across a set of layers of the same type.
+We first present ablations that isolate key design choices (dictionary capacity allocation, data-awareness, packing/quantization, and solver variants), and then report main results for both regimes.
 4.1 Experimental Setup
-We evaluate our method in both per-layer and cross-layer settings.
-For per-layer evaluations, we consider LLaMA-3.2 1B, Qwen-3 0.6B, LLaMA-3 8B, and Qwen-3 8B.
-For the cross-layer study, we follow the Basis Sharing Wang et al.
-(2024) setup and conduct experiments on LLaMA-2 7B.
-All models are evaluated in a zero-shot setting on PIQA Bisk et al.
-(2019), HellaSwag Zellers et al.
-(2019), OpenAI LAMBADA Paperno et al.
-(2016), ARC-easy and ARC-challenge Clark et al.
-(2018), SciQ Welbl et al.
-(2017), Race Lai et al.
-(2017), and MMLU Hendrycks et al.
-(2021), while perplexity is additionally reported on WikiText Merity et al.
-(2016) and LAMBADA-OpenAI.
-Compression is applied with target ratios of 0.2-0.5 with 0.1 step.
-For methods requiring calibration data, we randomly sample 256 sequences from the RefinedWeb dataset Penedo et al.
-(2023), each consisting of 1024 tokens.
-Unless specified otherwise, we compress all dense linear projections in self-attention (Q/K/V/O) and feed-forward network (gate/up/down).
-Embeddings and lm_head are left intact.
-For low-rank methods the rank is uniquely defined with $\gamma^{LR}$ and we floored it to the nearest integer.
-For CoSpaDi we fixed $\rho=2$ ($k/s$ ratio) for all experiments, so $k$ and $s$ were obtained according to Eq.
-(10) and then were floored to the nearest integers.
-In CoSpaDi we employ K-SVD using power iterations instead of full-svd for the dictionary updates.
-Specifically, we use 60 K-SVD iterations and 8 power iterations to ensure stable convergence.
-Further implementation details and convergence analysis are provided in Appendix A.8.
-4.2 Ablation Study
-Influence of $k/s$-ratio.
-In the proposed CoSpaDi we can redistribute capacity across both $k$ and $s$ while preserving predefined compression ratio by varying the $\rho$ value ($k/s$).
-We verified its influence on the Llama 3.2 1B and report the metrics in Fig.
-2.
-Figure 2: Dual-axis plot showing average accuracy (  solid lines, left axis) and perplexity (--- dashed lines, right axis, logarithmic scale with inverted direction) as functions of $\rho$ for Llama3.2-1B under three compression levels: 0.2, 0.3 and 0.4.
-Perplexity decreases upward due to axis inversion.
-CR Bitwidth Avg. Acc. PPL
-0.1686 bFP16 0.6198 1.94E+01
-0.1843 bFP15 0.6195 1.95E+01
-0.2001 bFP14 0.6176 1.97E+01
-0.2726 bFP16 0.5408 4.34E+01
-0.2864 bFP15 0.5394 4.37E+01
-0.3002 bFP14 0.5373 4.46E+01
-0.3765 bFP16 0.4096 1.80E+02
-0.3883 bFP15 0.4086 1.82E+02
-0.4002 bFP14 0.4069 1.84E+02
-0.4804 bFP16 0.2846 8.23E+02
-0.4902 bFP15 0.2838 8.15E+02
-0.5001 bFP14 0.2837 8.10E+02
-Table 1: Results on truncation of bfloat16 mantissa bits of coefficient matrix with reported average accuracy (Avg.
-Acc.) and Wiki Text word perplexity (PPL) for Llama3-8B resulting in different compression ratios (CR).
-From the provided plots we can observe that depending on the compression ratio the optimal $ks$-ratio differs, thus, for simplicity in further experiments we select in all cases $\rho=2$.
-Data-Free and Data-Aware Scenarios We argue that our proposed CoSpaDi allows for a more flexible representation compared to low-rank approximation, regardless of whether data-free or data-aware scenarios are considered.
-To prove this claim we performed an ablation study on Llama3-1B for different compression levels.
-We selected SVD for the data-free scenario and SVD-LLM for the data-aware case, while varying the compression ratio from 0.2 to 0.5.
-In Table 4.2 we report the results only for 0.2-0.5 compression.
-These results clearly indicate that in both data-free and data-aware settings the SDL-based methods outperform low-rank baselines by a wide margin, while CoSpaDi outperforms all methods.
+Models and layers.
+We evaluate per-layer compression on LLaMA-3.2-1B, Qwen-3-0.6B, LLaMA-3-8B, Qwen-3-8B and Qwen-3-14B.
+For SVD-LLM we used the original code-base [^1] https://github.com/AIoT-MLSys-Lab/SVD-LLM with only first step on compression without extra finetuning.
+For grouped compression, we follow the Basis Sharing [^2] https://github.com/TUDa-HWAI/Basis_Sharing protocol and report results on LLaMA-2-7B.
+Unless stated otherwise, we compress all dense linear projections in self-attention (Q/K/V/O) and gated MLP (up/down/gate), while leaving embeddings and the LM head intact.
+Calibration data and metrics.
+For calibration we randomly sample 256 sequences of length 1024 from RefinedWeb (Penedo et al., 2023).
+We report standard zero-shot accuracy (normalized when available) on a suite of benchmarks (PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), OpenAI LAMBADA (Paperno et al., 2016), ARC-easy/ARC-challenge (Clark et al., 2018), SciQ (Welbl et al., 2017), Race (Lai et al., 2017), MMLU (Hendrycks et al., 2021)) and perplexity on WikiText (Merity et al., 2017) and OpenAI LAMBADA.
+We used lm-evaluation-harness 0.4.8 (Gao et al., 2024) to ensure reproducibility.
+Compression parameterization.
+For low-rank baselines, the target compression ratio (CR) uniquely determines the rank $r$.
+For CoSpaDi, CR is achieved by choosing the dictionary size $k$ and sparsity $s$ (nonzeros per column), with $\rho\coloneqq k/s$ controlling how capacity is split between breadth (more atoms) and depth (more active atoms per column).
+Baselines.
+In the per-layer setting we compare primarily to SVD-LLM (data-aware) (Wang et al., 2025c) and, where relevant, to plain truncated SVD (data-free).
+We also include comparisons to training-free structured pruning baselines.
+We intentionally omit SVD-LLM v2 (Wang et al., 2025b) and Dobi-SVD (Qinsi et al., 2025) from the main baseline set for scope and fairness.
+Both methods emphasize dynamic (layer-wise) compression allocation, which is largely orthogonal to our contribution.
+Such allocation strategies are not specific to SVD and could be transferred to CoSpaDi by choosing $(k_{\ell},s_{\ell})$ per layer under a global budget.
+In addition, SVD-LLM v2 does not provide an official implementation, making it difficult to reproduce calibration and allocation details reliably and to apply them consistently in a dictionary-learning setting.
+Dobi-SVD further departs from our training-free protocol by requiring backpropagation to optimize the allocation.
+Its reported gains are also coupled to a quantization-specific “remapping” representation that changes the effective storage accounting.
+In particular, at moderate compression levels (e.g., CR <0.5<0.5), packing two 16-bit values into a single unit can lead to an effective overparameterization relative to standard bf16 storage.
+For these reasons, we focus on reproducible training-free baselines with standard storage conventions, and view dynamic allocation and quantization-specific representations as complementary extensions.
+4.2 Ablation Studies
+4.2.1 Capacity Allocation: the $k/s$ Ratio
+At a fixed compression ratio (CR), CoSpaDi admits a family of factorizations parameterized by $\rho\coloneqq k/s$.
+Smaller $\rho$ increases per-column expressiveness (larger $s$) but reduces the dictionary size $k$; larger $\rho$ increases $k$ but restricts each column to fewer active atoms.
+We perform a coarse sweep over $\rho\in[1;5]$.
+Figure 2 (LLaMA-3.2-1B) shows that $\rho=2$ consistently provides the best trade-off across all considered CRs in terms of both average accuracy and perplexity.
+We therefore fix $\rho=2$ in all subsequent experiments for simplicity and comparability across models.
+Notably, $\rho=1$ corresponds to the degenerate regime $k=s$, where each column may activate all atoms.
+In this case, the sparsity constraint becomes inactive and CoSpaDi reduces to a dense two-factor approximation, closely mirroring the “basis times coefficients” view underlying truncated SVD/PCA.
+This emphasizes that CoSpaDi can be seen as a strict generalization: choosing $\rho>1$ enforces sparsity and yields a union-of-subspaces model.
+Figure 2: Dual-axis plot showing average accuracy (  solid lines, left axis) and WikiText perplexity (--- dashed lines, right axis, inverted logarithmic scale) as functions of $\rho$ for Llama3.2-1B under three compression levels: 0.2, 0.3 and 0.4.
+Figure 3: Dual-axis plot showing relative average accuracy (  solid lines, left axis) and relative WikiText perplexity (--- dashed lines, right axis) as functions of $\mathbf{S}$ bitwidth for Llama3-8B under three compression levels: 0.2, 0.3 and 0.4.
+4.2.2 Data-Free vs. Data-Aware Compression
+A core motivation for dictionary learning is that its union-of-subspaces form can be more expressive than a single low-dimensional subspace regardless of whether calibration is used.
+We therefore compare: (i) data-free truncated SVD vs.
+CoSpaDi (dictionary learning in weight space, without whitening), and (ii) data-aware SVD-LLM vs.
+CoSpaDi (activation-aware via whitening).
+Table 2 summarizes results on LLaMA-3.2-1B across CR 0.2–0.4.
+In both regimes, the dictionary-based factorization is consistently stronger than the corresponding low-rank baseline, and activation-aware CoSpaDi yields the best overall trade-off.
 Table 1: SDL-based methods comparison vs low-rank counterparts in data-free and data-aware scenarios on Llama3.2-1B at different compression ratios (CR).
@@
 Method Data-Aware CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
+Llama3.2 1B 74.5 63.7 63.0 60.5 36.2 88.3 37.8 37.0 57.6 11.6E+00 5.7E+00
+SVD
+52.0 25.7 0.0 24.7 22.1 20.0 21.4 25.4 23.9 2.9E+06 4.6E+06
+$\text{CoSpaDi}^{\dagger}$
+51.7 26.4 0.0 25.4 25.3 21.0 21.8 24.1 24.5 3.3E+05 2.2E+06
+SVD-LLM
+62.1 36.4 24.4 36.0 25.1 64.9 29.0 23.0 37.6 1.7E+02 1.7E+02
+CoSpaDi
+0.2 66.1 42.9 38.4 39.9 26.0 71.6 31.7 24.8 42.7 6.4E+01 3.5E+01
+SVD
+52.3 25.6 0.0 24.2 23.5 19.5 21.9 27.0 24.3 1.1E+06 3.9E+06
+$\text{CoSpaDi}^{\dagger}$
+50.5 26.3 0.0 24.5 26.1 21.8 21.5 25.5 24.5 2.1E+05 4.3E+06
+SVD-LLM
+55.7 30.1 9.1 30.5 21.5 45.9 25.8 23.2 30.2 5.9E+02 2.5E+03
+CoSpaDi
+0.3 56.9 32.4 18.2 31.9 22.1 56.7 28.0 23.1 33.7 2.9E+02 6.6E+02
+SVD
+52.8 25.9 0.0 23.9 21.3 19.9 22.2 26.9 24.1 1.2E+06 4.2E+06
+$\text{CoSpaDi}^{\dagger}$
+51.0 26.3 0.0 25.5 26.9 21.3 21.2 25.5 24.7 3.1E+06 3.7E+07
+SVD-LLM
+51.8 27.3 1.3 26.9 22.9 32.3 24.4 23.0 26.2 1.6E+03 3.3E+04
+CoSpaDi
+0.4 53.5 28.2 3.8 27.8 23.0 36.9 24.0 23.1 27.5 8.0E+02 9.2E+03
+4.2.3 Dictionary Learning Solver Variants: Exact K-SVD vs. Power-K-SVD vs. MOD
+The main bottleneck in the alternating minimization process is the dictionary update using K-SVD as it updates a single atom at a time.
+To reduce its computational cost we investigate the usage of truncated PCA and power iterations for solving rank-1 problems as well as the well-known Method of Optimal Directions (MOD) which updates in a single step the entire dictionary instead of individual atoms sequentially.
+Table 2 reports both quality (Avg.
+accuracy, WikiText perplexity) and wall-clock compression time at a representative 0.2 CR on LLaMA-3.2-1B.Wall-clock is measured for compressing all targeted projection matrices with identical hardware, batching, and stopping criteria.
+This comparison isolates the practical speed–quality trade-off of the solver choice while keeping the factorization form $(k,s)$ fixed.
+For our further experiments we utilized K-SVD using power-iterations as it achieves good performance at a reasonable wall-clock time for compression.
+We provide a more thorough ablation study on the number of power iterations and of the alternating steps in Appendix A.7
+Table 2: Solver comparison for dictionary update on LLaMA-3.2-1B at 0.2 compression ratio (CR) and fixed $\rho=2$, 60 alternating minimization iterations.
+We report accuracies, perplexity as well as wall-clock compression time in minutes using a single A100.
+Best results are highlighted with bold.
 […]
-SVD x 0.5201 0.2566 0.0003 0.2471 0.2210 0.2000 0.2144 0.2537 0.2391 2.93E+06 4.56E+06
-$\text{CoSpaDi}^{\dagger}$ x 0.5174 0.2635 0.0002 0.2538 0.2534 0.2100 0.2182 0.2413 0.2447 3.33E+05 2.16E+06
-SVD-LLM v 0.6213 0.3643 0.2443 0.3603 0.2509 0.6490 0.2900 0.2298 0.3762 1.69E+02 1.70E+02
-CoSpaDi v 0.2 0.6605 0.4288 0.3839 0.3994 0.2602 0.7160 0.3167 0.2483 0.4267 6.37E+01 3.51E+01
-SVD x 0.5234 0.2564 0.0002 0.2416 0.2346 0.1950 0.2191 0.2698 0.2425 1.09E+06 3.92E+06
-$\text{CoSpaDi}^{\dagger}$ x 0.5049 0.2626 0.0002 0.2449 0.2611 0.2180 0.2153 0.2552 0.2453 2.06E+05 4.34E+06
-SVD-LLM v 0.5565 0.3006 0.0910 0.3047 0.2150 0.4590 0.2583 0.2323 0.3022 5.90E+02 2.47E+03
-CoSpaDi v 0.3 0.5691 0.3241 0.1822 0.3194 0.2210 0.5670 0.2804 0.2308 0.3368 2.89E+02 6.59E+02
-SVD x 0.5277 0.2590 0.0001 0.2386 0.2133 0.1990 0.2220 0.2693 0.2411 1.23E+06 4.17E+06
-$\text{CoSpaDi}^{\dagger}$ x 0.5103 0.2626 0.0001 0.2546 0.2688 0.2130 0.2115 0.2549 0.2470 3.05E+06 3.67E+07
-SVD-LLM v 0.5180 0.2725 0.0126 0.2685 0.2287 0.3230 0.2440 0.2297 0.2621 1.58E+03 3.30E+04
-CoSpaDi v 0.4 0.5348 0.2824 0.0380 0.2778 0.2295 0.3690 0.2402 0.2310 0.2753 7.96E+02 9.23E+03
-SVD x 0.5392 0.2566 0.0001 0.2441 0.2039 0.2040 0.2077 0.2690 0.2406 1.22E+06 1.88E+07
-$\text{CoSpaDi}^{\dagger}$ x 0.5120 0.2607 0.0001 0.2572 0.2654 0.2280 0.2211 0.2704 0.2519 6.09E+05 1.08E+07
-SVD-LLM v 0.5109 0.2663 0.0004 0.2609 0.2594 0.2610 0.2392 0.2303 0.2536 3.13E+03 1.03E+05
-CoSpaDi v 0.5 0.5169 0.2703 0.0031 0.2626 0.2398 0.2950 0.2421 0.2326 0.2578 1.80E+03 7.26E+04
-Quantization of Sparse Coefficient Matrix We investigated the quantization of the coefficient matrix $\mathbf{S}$ in the post-training regime to leverage the memory add-on due to the requirement of storing the indices of the nonzero values.
-We utilized a naive mantissa bit truncation of the original bfloat16 coefficient values.
-In Table 1 we report how truncation affects the performance at different compression levels.
-According to the obtained results, truncation of 2 bits leads to negligible drop.
-Thus, in all our experiments CoSpaDi consists of bfloat16 dictionary and coefficient matrices but is evaluated with the truncated version of the sparse coefficient matrix using 14-bits.
+Method time PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
+Llama3.2 1B 74.5 63.7 63.0 60.5 36.2 88.3 37.8 37.0 57.6 11.6E+00 5.7E+00
+MOD 222.2 62.9 39.7 30.4 36.3 25.7 67.6 27.9 23.0 39.2 1.1E+02 8.7E+01
+K-SVD (PCA) 902.1 64.4 39.9 33.7 40.5 25.9 72.3 31.6 24.4 41.6 9.6E+01 6.2E+01
+K-SVD (power) 646.8 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
+4.2.4 Quantizing Coefficients and Removing the Mask Term
+Sparse factorization introduces metadata overhead for storing nonzero locations.
+Our default implementation uses a packed binary mask (one bit per entry) plus bf16 values for the $sd_{2}$ nonzeros.
+We additionally study post-training coefficient quantization by truncating bf16 mantissa bits.
+Plots on Figure 3 shows that truncating 2 mantissa bits yields negligible degradation across tested CRs.
+Beyond robustness, coefficient quantization enables a simplified storage accounting in which locations can be treated as fixed by construction (e.g., using a deterministic pattern per column or a reproducible pseudo-random seed), removing the explicit mask term from the compression ratio.
+Under this view, the effective storage becomes
+$\hat{\gamma}^{\mathrm{SD}}=1-\frac{d_{1}k+sd_{2}}{d_{1}d_{2}}.$ (14)
+Unless stated otherwise, we evaluate CoSpaDi with bf16 dictionaries and truncated (14-bit) coefficients, thus providing CR according to Equation 14.
 4.3 Main Results
-Per-Layer Scenario:
-In this section we apply CoSpaDi for each projection layer independently and compare it against current sota training-free low-rank method – SVD-LLM Wang et al. (2025a).
-(a) Accuracy & Perplexity for LLaMA3.2-1B
-(b) Accuracy & Perplexity for Qwen3-0.6B
-Figure 3: Average benchmark accuracy and WikiText perplexity for (a) LLaMA-3.2-1B and (b) Qwen-3 0.6B using SVD-LLM and CoSpaDiwith respect to compression ratio.
-From the provided plots CoSpaDi significantly outperforms SVD-LLM in both avg.
-accuracy and perplexity for small Llama3 and Qwen3 models.
-We further investigate how our method scales to larger models of the same families with 8B parameters.
-The related results are provided in Table 4.3.
-We have also performed comparisons on the same models and compression levels against other training-free compression strategies.
-In particular we compare CoSpaDi against ASVD Yuan et al.
-(2023), ReplaceMe Shopkhoev et al.
-(2025) and LLMPruner Ma et al.
-(2023), with the last two being sota pruning methods.
-These results are provided in Appendix A.6
-Table 3: Performance comparison of CoSpaDi vs sota SVD-LLM on Llama3-8B and Qwen3-8B at different compression levels on different benchmarks. Best results are highlighted with bold.
+4.3.1 Per-Layer Compression
+We first compress each projection matrix independently and compare CoSpaDi to SVD-LLM over a range of CRs.
+Figure 4 summarizes the accuracy–compression and perplexity–compression curves for small models (LLaMA-3.2-1B and Qwen-3-0.6B).
+For larger models, Table 3 reports results on LLaMA-3-8B, Qwen-3-8B and Qwen-3-14B at CR 0.2–0.4.
+Across model families and budgets, CoSpaDi consistently improves both average benchmark accuracy and perplexity at matched CR, indicating that the union-of-subspaces factorization better preserves task-relevant directions than a single low-rank basis.
+We provide additional benchmarks in the Appendix A.6.
+Figure 4: Average benchmark accuracy and WikiText perplexity for LLaMA-3.2-1B and Qwen-3 0.6B using SVD-LLM and CoSpaDi with respect to compression ratio.
+Table 3: Performance comparison of CoSpaDi vs sota SVD-LLM on Llama3-8B, Qwen3-8B and Qwen3-14B at different compression levels on different benchmarks. Best results are highlighted with bold.
 […]
 Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
+Llama3 8B 80.7 79.1 75.6 77.7 53.5 93.9 40.3 62.2 70.4 7.3E+00 3.1E+00
+SVD-LLM 71.1 58.4 59.3 55.5 34.0 86.4 35.5 32.6 54.1 4.1E+01 1.1E+01
+CoSpaDi 0.2 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
+SVD-LLM 65.8 46.4 38.1 41.9 27.7 70.0 31.8 27.2 43.6 1.5E+02 6.1E+01
+CoSpaDi 0.3 70.5 56.2 61.3 54.2 33.5 85.7 36.2 32.2 53.7 4.5E+01 9.2E+00
+SVD-LLM 60.3 34.5 11.4 32.4 24.5 44.2 25.7 23.1 32.0 5.5E+02 1.3E+03
+CoSpaDi 0.4 63.7 41.4 30.3 39.1 26.6 68.5 30.5 25.4 40.7 1.8E+02 1.2E+02
+Qwen3 8B 77.7 74.9 64.1 80.7 56.7 95.7 40.9 73.0 70.5 1.2E+01 4.6E+00
+SVD-LLM 73.8 63.9 62.2 68.7 45.7 90.1 40.5 54.7 62.5 2.1E+01 6.4E+00
+CoSpaDi 0.2 76.5 68.0 65.6 72.2 48.9 93.2 40.7 60.8 65.7 1.8E+01 4.9E+00
+SVD-LLM 70.4 55.2 53.8 59.3 37.1 87.2 38.4 44.8 55.8 2.7E+01 1.1E+01
+CoSpaDi 0.3 72.4 60.5 62.6 63.9 41.2 88.4 39.5 51.3 59.97 2.3E+01 6.3E+00
+SVD-LLM 66.3 44.6 37.9 45.0 28.1 77.3 35.3 29.1 45.4 4.3E+01 3.6E+01
+CoSpaDi 0.4 68.9 49.0 49.9 49.4 29.9 82.0 36.8 36.6 50.3 3.6E+01 1.5E+01
+Qwen3 14B 79.9 78.9 67.9 82.8 60.2 96.5 43.3 77.2 73.3 1.1E+01 3.7E+00
+SVD-LLM 76.2 67.6 69.1 69.8 46.8 91.0 42.9 62.5 65.8 1.8E+01 4.1E+00
+CoSpaDi 0.2 77.3 71.7 72.0 73.3 51.3 91.9 43.1 65.8 68.3 1.6E+01 3.4E+00
+SVD-LLM 72.0 59.7 61.5 62.6 39.3 88.9 41.5 52.0 59.7 2.3E+01 6.5E+00
+CoSpaDi 0.3 74.6 64.2 69.0 65.3 43.6 90.7 42.5 58.0 63.5 2.0E+01 4.2E+00
+SVD-LLM 67.4 48.7 46.0 48.5 30.7 80.6 36.7 36.8 49.4 3.5E+01 1.8E+01
+CoSpaDi 0.4 69.5 52.3 57.3 53.2 33.0 84.4 37.5 43.6 53.9 3.3E+01 8.7E+00
+4.3.2 Comparison with Pruning
+We additionally compare CoSpaDi to recent structured pruning approaches on the same models and compression ratios.
+Specifically, we evaluate against ReplaceMe (Shopkhoev et al., 2025) and LLM-Pruner (Ma et al., 2023), which are representative state-of-the-art methods that reduce parameters by removing or replacing structured components (e.g., blocks/layers) rather than individual weights.
+Results are reported in Table 4.
+We do not include unstructured pruning baselines in this comparison because their practical memory savings depend on sparse storage formats and hardware support; at moderate sparsity levels (e.g., $\approx$50%), the overhead of storing indices and irregular access patterns often diminishes or eliminates effective compression and speedups in standard dense inference pipelines.
+Table 4: Comparison of the proposed CoSpaDi method with state-of-the-art structured pruning methods ReplaceMe Shopkhoev et al.
+(2025) and LLM-Pruner Ma et al.
+(2023) on Llama3 8B under different compression ratios.
+We report accuracy on different benchmarks as well as its average and perplexity.
+Best results are highlighted with bold
 […]
-SVD-LLM 0.7106 0.5837 0.5925 0.5551 0.3396 0.8640 0.3550 0.3262 0.5408 4.07E+01 1.09E+01
-CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
-SVD-LLM 0.6578 0.4640 0.3806 0.4188 0.2765 0.7000 0.3177 0.2715 0.4358 1.47E+02 6.09E+01
-CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
-SVD-LLM 0.6028 0.3450 0.1143 0.3237 0.2449 0.4420 0.2574 0.2305 0.3201 5.49E+02 1.30E+03
-CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
-SVD-LLM 0.5381 0.2790 0.0041 0.2731 0.2389 0.2660 0.2230 0.2297 0.2565 1.37E+03 2.80E+04
-CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
+Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
+Llama3 8B 80.7 79.1 75.6 77.7 53.5 93.9 40.3 62.2 70.4 7.3E+00 3.1E+00
+ReplaceMe 0.22 73.1 65.7 42.1 65.9 43.7 86.4 35.4 51.7 58.0 3.4E+01 2.0E+01
+LLM-Pruner 75.5 67.5 51.0 62.1 36.6 87.8 35.1 25.0 55.1 1.6E+01 1.1E+01
+CoSpaDi 0.2 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
+ReplaceMe 0.31 66.6 53.8 24.0 50.7 37.9 77.3 34.0 30.6 46.9 6.7E+01 1.3E+02
+LLM-Pruner 67.3 45.1 20.9 45.4 28.8 63.4 30.1 22.9 40.5 3.8E+01 2.2E+02
+CoSpaDi 0.3 70.5 56.2 61.3 54.2 33.5 85.7 36.2 32.2 53.7 4.5E+01 9.2E+00
+ReplaceMe 0.41 61.7 44.3 9.8 37.4 27.5 60.4 31.6 26.4 37.4 2.3E+02 1.8E+03
+LLM-Pruner 50.3 25.8 1.5 26.4 25.8 28.1 21.8 23.2 25.4 $\infty$ 5.7E+05
+CoSpaDi 0.4 63.7 41.4 30.3 39.1 26.6 68.5 30.5 25.4 40.7 1.8E+02 1.2E+02
+4.3.3 Grouped (Cross-Layer) Dictionary Sharing
+Next, we evaluate grouped compression where a single dictionary is shared across multiple layers of the same projection type.
+For a fair and controlled comparison, we adopt the exact layer grouping and evaluation protocol of Basis Sharing, i.e., we share parameters across the same layer sets and report results under the same compression budgets.
+Table 5 reports results on LLaMA-2-7B across CR 0.2–0.4.
+We additionally included per-layer counterparts to show the benefits of grouping strategy and exploiting inter-layer redundancies.
+At all budgets, CoSpaDi substantially outperforms Basis Sharing as well as per-layer baselines, suggesting that cross-layer sharing is most effective when paired with sparse codes that allow each column (and each layer) to select its own subset of atoms.
+We note that CoSpaDi does not rely on this particular grouping and could benefit from more adaptive sharing strategies (e.g., data-driven grouping or partially shared dictionaries), which we leave for future work.
+Table 5: Performance comparison of CoSpaDi vs Basis Sharing (Wang et al., 2025a) and per-layer counterparts on Llama2-7B under different compression levels on various benchmarks.
+Best results are highlighted with bold.
 […]
-SVD-LLM 0.7383 0.6391 0.6218 0.6869 0.4573 0.9010 0.4048 0.5473 0.6246 2.06E+01 6.40E+00
-CoSpaDi 0.2 0.7650 0.6804 0.6559 0.7218 0.4889 0.9320 0.4067 0.6075 0.6573 1.81E+01 4.88E+00
-SVD-LLM 0.7035 0.5522 0.5377 0.5926 0.3712 0.8720 0.3837 0.4482 0.5576 2.74E+01 1.07E+01
-CoSpaDi 0.3 0.7242 0.6050 0.6255 0.6385 0.4121 0.8840 0.3952 0.5127 0.5997 2.29E+01 6.29E+00
-SVD-LLM 0.6627 0.4458 0.3790 0.4503 0.2807 0.7730 0.3531 0.2901 0.4544 4.30E+01 3.59E+01
-CoSpaDi 0.4 0.6888 0.4898 0.4991 0.4941 0.2986 0.8200 0.3675 0.3660 0.5030 3.59E+01 1.47E+01
-SVD-LLM 0.6023 0.3543 0.1956 0.3363 0.2278 0.6670 0.2909 0.2305 0.3631 8.72E+01 2.69E+02
-CoSpaDi 0.5 0.6159 0.3798 0.2791 0.3565 0.2295 0.6860 0.3167 0.2305 0.3868 6.82E+01 1.10E+02
-Cross-Layer Scenario
-In this section, we validate CoSpaDi on the cross-layer compression scenario, where a common dictionary is shared across different layers.
-To ensure a fair comparison, we adopt the identical layer selection and grouping procedure introduced in Basis Sharing Wang et al.
-(2024) (for a detailed description we refer to Appendix A.7) and compare the performance of the two methods on the Llama2-7B model Touvron et al.
-(2023b).
-As shown in Table LABEL:tab:main_results_group, CoSpaDi consistently outperforms Basis Sharing by a large margin across all benchmarks and compression rates (0.2–0.5).
-This performance gap suggests that CoSpaDi’s sparsity-aware, dictionary-based decomposition better preserves task-critical features under aggressive compression.
-We believe that a more sophisticated layer grouping strategy, based on feature map similarity or end-to-end differentiable search, could work better with CoSpaDi and further widen this performance margin, but we leave this as a future research direction.
-We provide details for $k$ (dictionary size) and $s$ (sparsity) parameters in the Appendix A.7.
+Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
+Llama2 7B 78.0 57.1 73.7 76.1 43.3 93.9 39.2 40.9 62.8 8.7 3.4
+SVD-LLM 73.3 60.4 64.0 55.1 32.3 79.7 36.6 25.5 53.4 20.2 6.3
+CoSpaDi (per-layer) 74.3 64.8 70.2 61.6 35.3 85.8 39.0 28.2 57.4 15.8 4.5
+Basis Sharing 70.1 43.4 63.0 66.5 33.7 91.0 34.9 24.8 53.4 15.2 7.2
+CoSpaDi (grouped) 0.2 74.6 66.3 71.3 67.4 39.3 88.8 38.3 27.2 59.2 11.7 4.3
+SVD-LLM 68.4 52.2 51.5 46.8 27.8 75.9 34.7 23.5 47.6 34.0 13.5
+CoSpaDi (per-layer) 71.2 57.4 62.7 53.5 30.6 81.6 36.8 26.9 52.6 23.6 6.9
+Basis Sharing 65.6 37.6 53.5 58.8 27.3 86.8 32.6 23.2 48.2 22.2 13.6
+CoSpaDi (grouped) 0.3 70.8 58.5 64.5 61.9 34.3 87.6 35.9 24.0 54.7 15.4 6.3
+SVD-LLM 63.2 42.7 33.1 37.7 24.6 70.2 32.2 23.0 40.8 86.5 67.1
+CoSpaDi (per-layer) 66.2 48.4 47.2 44.3 26.2 76.7 33.4 23.4 45.7 49.2 20.9
+Basis Sharing 60.6 33.1 41.1 49.5 23.6 83.0 29.7 23.1 42.95 39.5 37.7
+CoSpaDi (grouped) 0.4 64.3 48.0 51.8 52.3 29.9 81.5 33.2 23.0 48.0 25.2 14.6
+5 Open Issues and Future Work
+Solver efficiency and scalability.
+CoSpaDi uses alternating minimization with OMP for sparse coding and K-SVD-style sequential atom updates.
+This is simple and effective but can be expensive due to repeated sparse coding and per-atom updates, especially for large matrices and aggressive compression.
+Faster variants (e.g., more parallel updates, better batching, or alternative sparse coding) are promising; our ablations already show that approximate rank-1 updates (e.g., power iterations) offer a favorable accuracy–time trade-off.
+Hyperparameter selection and dynamic budget allocation.
+We parameterize CoSpaDi via $(k,s)$ and $\rho=k/s$, but fix $\rho$ and use uniform budgets for simplicity.
+Prior low-rank work shows that layer-/type-wise sensitivity can improve quality at a fixed global budget.
+Adapting dynamic allocation to dictionary learning–e.g., choosing $(k_{\ell},s_{\ell})$ per layer, projection type, or group based on functional sensitivity–is a promising direction.
+Grouped sharing design choices.
+Grouped performance depends on how layers are clustered and which projections share a dictionary.
+We follow a simple grouping aligned with prior cross-layer sharing, while more adaptive grouping based on activation/weight similarity could yield additional gains.
+Practical speedups depend on kernel support.
+CoSpaDi yields structured sparsity exploitable by sparse–dense computation, but end-to-end speedups depend on sparse kernel efficiency, indexing overhead, and reuse of intermediate products.
+Encouragingly, recent work reports strong acceleration for dense–sparse matmul even at modest sparsity, narrowing the gap between parameter reduction and wall-clock benefits (Macko & Boža, 2025).
+Realizing consistent speedups will require integration into optimized inference runtimes and hardware-aware kernels for the resulting computation patterns.
+Beyond the current instantiation.
+Our goal is to introduce activation-aware sparse dictionary learning as a new post-training compression paradigm beyond single-subspace (low-rank) approximations.
+Accordingly, we use a direct optimization pipeline.
+Many advances from low-rank compression–richer calibration objectives, better sensitivity measures, and refined allocation strategies–could be transferred to this setting without changing the core factorization.
+6 Conclusion
+We presented CoSpaDi, a training-free LLM compression framework based on sparse dictionary learning.
+CoSpaDi factorizes projection matrices into a dense dictionary and column-sparse codes, moving from a single shared low-dimensional subspace to a union-of-subspaces representation.
+Using a small calibration set, an activation-derived Gram orthonormalization converts functional output reconstruction into a tractable dictionary learning objective solved by alternating sparse coding and dictionary updates.
+Across multiple Llama and Qwen models, CoSpaDi improves the accuracy–compression and perplexity–compression trade-offs over strong SVD-based baselines in both per-layer and grouped (shared-dictionary) settings at 20–40% compression ratios, and composes naturally with post-training coefficient quantization.
+More broadly, our results highlight dictionary learning as a flexible paradigm for post-training compression and suggest that many advances from low-rank methods (e.g., dynamic allocation and improved calibration objectives) can be transferred to this richer factorization family.
 Appendix A Appendix
 A.1 Derivation of the optimal pair of basis and coefficient matrices
-We are seeking the optimal pair $\left(\mathbf{B}^{\star},\mathbf{C}^{\star}\right)$ that minimizes the constrained problem in Eq.
-(3).
-First, we rewrite the objective $\mathcal{J}$ in its equivalent form:
+Let us consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$.
+We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$.
+The optimal basis and coefficients can be found by minimizing the total approximation error:
+$\mathcal{J}=\sum\limits_{j=1}^{d_{2}}\left\|\mathbf{w}_{j}-\sum\limits_{i=1}^{r}c_{i,j}\bm{b}_{i}\right\|_{2}^{2}\equiv\left\|\mathbf{W}-\mathbf{B}\mathbf{C}\right\|_{F}^{2},$ (15)
+subject to the orthogonality constraint $\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}$, where $\mathbf{B}=\begin{bmatrix}\bm{b}_{1},\ldots,\bm{b}_{r}\end{bmatrix}\in\mathbb{R}^{d_{1}\times r}$ is the basis matrix, $\mathbf{C}\in\mathbb{R}^{r\times d_{2}}$ is the coefficient matrix with entries $c_{i,j}$, and $\mathbf{I}\in\mathbb{R}^{r\times r}$ is the identity matrix.
+To solve this problem, we first reformulate the objective $\mathcal{J}$ into an equivalent form:
 $\mathcal{J}=\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}\right)-2\operatorname*{tr}\left(\mathbf{W}^{\mathsf{T}}\mathbf{B}\mathbf{C}\right)+\operatorname*{tr}\left(\mathbf{C}^{\mathsf{T}}\mathbf{B}^{\mathsf{T}}\mathbf{B}\mathbf{C}\right).$ (16)
@@
 This yields the compression ratio in Appendix A.4 and Eq.
-9.
+12.
 (5) Inference.
@@
 $O_{SD,sparse-reuse}=Nd_{1}K_{active}+Nsd_{2},\qquad s\leq K_{active}\leq\min(K,sd_{2}).$ (30)
-With proposed truncation of 2-bits in mantissa we can omit term for storing indices of nonzero elements in $\mathbf{S}$ resulting in corrected compression ratio: ^γ^SD = 1 - d1k + s d2d1d2 The rank for the low-rank decomposition could be estimated from the compression ratio with the following equation: r=(1 - γLR) d1d2d1+ d2 For sparse dictionary based method we defined $\rho=k/s$ and, thus we can estimate both $k$ and $s$ in the following way: k=(1-^γSD) d1d2d1+d2ρ,  s=kρ.
-Under the same compression ratio for both SVD and CoSpaDi we obtain exactly the same complexity: O_LR=N r (d_1 + d_2) = N (1 - γ^LR) d_1 d_2 O_SD=N d_1 k + N s d_2 = N(d_1 k + d_2 kρ) = N k (d_1 + d2ρ) = N (1-^γSD) d1d2d1+d2ρ(d_1 + d2ρ)=N(1-^γ^SD) d_1d_2
+With proposed truncation of 2-bits in mantissa we can omit term for storing indices of nonzero elements in $\mathbf{S}$ resulting in corrected compression ratio:
+$\hat{\gamma}^{SD}=1-\frac{d_{1}k+sd_{2}}{d_{1}d_{2}}$
+The rank for the low-rank decomposition could be estimated from the compression ratio with the following equation:
+$r=\frac{(1-\gamma^{LR})d_{1}d_{2}}{d_{1}+d_{2}}$
+For sparse dictionary based method we defined $\rho=k/s$ and, thus we can estimate both $k$ and $s$ in the following way:
+$k=\frac{(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}\,},\qquad s=\tfrac{k}{\rho}.$
+Under the same compression ratio for both SVD and CoSpaDi we obtain exactly the same complexity:
+$O_{LR}=Nr(d_{1}+d_{2})=N(1-\gamma^{LR})d_{1}d_{2}$
+$O_{SD}=Nd_{1}k+Nsd_{2}=N(d_{1}k+d_{2}\frac{k}{\rho})=Nk(d_{1}+\frac{d_{2}}{\rho})=N\frac{(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}\,}(d_{1}+\frac{d_{2}}{\rho})=N(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}$
 $O_{LR}=O_{SD}=N(1-\gamma)d_{1}d_{2}.$ (31)
 While, theoretical complexity is the same, in practice inference time depends on the sparsity structure ($K_{\text{active}}$), indexing overhead, and kernel efficiency, as can be observed from Figs.
-4, 5 and   6.
+5, 6 and   7.
 Figure 5: Inference time for different projection layers of Llama3.2 1B for different compression ratios and $k/s$ ratios on A100
@@
 Figure 7: Inference time for different projection layers of Qwen3 0.6B for different compression ratios and $k/s$ ratios on A100
-A.6 Comparison with Other Training-Free Methods
-We also evaluate the performance of the proposed CoSpaDi relative to other training free methods, particularly structural pruning ones and ASVD Yuan et al.
-(2023) which is another data-aware SVD-based method.
-The results are provided in Table A.6.
-Across LLaMA-3 8B, our method consistently outperforms structural pruning baselines at moderate budgets.
-At $\text{CR}\!\approx\!0.2$, CoSpaDi attains the best average accuracy (0.618 vs.
-0.580 for ReplaceMe and 0.551 for LLM-Pruner) while also delivering the lowest perplexities, indicating balanced generative and discriminative quality.
-At $\text{CR}\!\approx\!0.3$, CoSpaDi widens the margin in average accuracy (0.537 vs.
-0.469/0.405) and avoids the severe perplexity blow-ups observed for pruning methods.
-Even under aggressive compression ($\text{CR}\!\approx\!0.4$–0.5), CoSpaDi remains competitive or superior on most tasks and preserves far more stable perplexity profiles, whereas ReplaceMe and LLM-Pruner exhibit task collapse (near-zero LAMBADA and extreme PPL spikes).
-Overall, the results support dictionary–sparsity as a more robust data-free alternative to structural pruning at matched or tighter budgets, especially when both accuracy and perplexity must be maintained.
-Table 5: Comparison of the proposed CoSpaDi method with other training-free methods including ASVD Yuan et al.
-(2023) and state-of-the-art structured pruning methods ReplaceMe Shopkhoev et al.
-(2025) and LLM-Pruner Ma et al.
-(2023) on Llama3 8B under different compression ratios.
-We report accuracy on different benchmarks as well as its average and perplexity.
-Best results are highlighted with bold
-[…]
-Method CR PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
-[…]
-ASVD 0.2 0.5577 0.2666 0.0070 0.3237 0.2150 0.3050 0.2182 0.2347 0.2660 9.38E+04 9.89E+05
-ReplaceMe 0.22 0.7307 0.6572 0.4211 0.6587 0.4369 0.8640 0.3541 0.5167 0.5799 3.37E+01 2.01E+01
-LLM-Pruner 0.7546 0.6746 0.5096 0.6208 0.3660 0.8780 0.3512 0.2495 0.5505 1.60E+01 1.05E+01
-CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
-ASVD 0.3 0.5359 0.2595 0.0000 0.2694 0.2142 0.2190 0.2115 0.2349 0.2431 3.43E+05 1.35E+07
-ReplaceMe 0.31 0.6659 0.5382 0.2401 0.5067 0.3788 0.7730 0.3397 0.3058 0.4685 6.67E+01 1.33E+02
-LLM-Pruner 0.6725 0.4509 0.2090 0.4541 0.2875 0.6340 0.3014 0.2292 0.4048 3.79E+01 2.21E+02
-CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
-ASVD 0.4 0.5365 0.2575 0.0000 0.2504 0.2287 0.2080 0.2201 0.2431 0.2430 2.05E+05 2.57E+07
-ReplaceMe 0.41 0.6170 0.4430 0.0976 0.3742 0.2747 0.6040 0.3158 0.2638 0.3738 2.30E+02 1.76E+03
-LLM-Pruner 0.5033 0.2580 0.0151 0.2643 0.2577 0.2810 0.2182 0.2324 0.2537 $\infty$ 5.67E+05
-CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
-ASVD 0.5196 0.2547 0.0000 0.2601 0.2261 0.1980 0.2134 0.2554 0.2409 1.72E+06 1.72E+07
-ReplaceMe 0.5724 0.3635 0.0367 0.2976 0.2543 0.3810 0.2785 0.2300 0.3017 6.99E+02 5.03E+04
-LLM-Pruner 0.5038 0.2623 0.0043 0.2630 0.2637 0.2560 0.2172 0.2371 0.2509 1.24E+03 1.23E+06
-CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
-A.7 Cross-Layer Scenario
-In Table 6, we report the dictionary size ($k$), sparsity level ($s$), and input/output weight dimensions for each linear transform under a fixed $k/s$ ratio.
-This includes the query, key, value, and output projections within the attention block, as well as the up, down, and gated projections in the gated MLP block.
-The dictionary and sparsity dimensions are adjusted according to the target compression rate.
-Table 6: Compression configurations for Llama2 7B weight matrices across varying compression rates (20%–50%).
-Each row specifies sparsity parameters ($k$, $s$, $k/s$ ratio) and weight dimensions ($d_{1}$, $d_{2}$) for different weight types (Query, Key, Value, Up, Gate, Down, Out), grouped by compression rate and group size (1 or 2).
-Weight Type Compression Rate Group size Dictionary, $k$ Sparsity, $s$ $k/s$ ratio d1 d2
-Query 20% 2 3276 1638 2 4096 8192
-Key 3276 1638 2 4096 8192
-Value 3276 1638 2 4096 8192
-Up 4776 2388 2 4096 22016
-Gate 4776 2388 2 4096 22016
-Down 1 2762 1381 2 11008 4096
-Out 2184 1092 2 4096 4096
-Query 30% 2 2866 1433 2 4096 8192
-Key 2866 1433 2 4096 8192
-Value 2866 1433 2 4096 8192
-Up 4178 2089 2 4096 22016
-Gate 4178 2089 2 4096 22016
-Down 1 2416 1208 2 11008 4096
-Out 1910 955 2 4096 4096
-Query 40% 2 2456 1228 2 4096 8192
-Key 2456 1228 2 4096 8192
-Value 2456 1228 2 4096 8192
-Up 3582 1791 2 4096 22016
-Gate 3582 1791 2 4096 22016
-Down 1 2072 1036 2 11008 4096
-Out 1638 819 2 4096 4096
-Query 50% 2 2048 1024 2 4096 8192
-Key 2048 1024 2 4096 8192
-Value 2048 1024 2 4096 8192
-Up 2984 1492 2 4096 22016
-Gate 2984 1492 2 4096 22016
-Down 1 1726 863 2 11008 4096
-Out 1364 682 2 4096 4096
-A.8 K-SVD and Power Iteration Analysis
+A.6 More Benchmarks
+We additionally evaluate CoSpaDi and SVD-LLM on the Qwen model family using a set of more recent reasoning-oriented benchmarks.
+Table 6 reports results for Qwen3-8B and Qwen3-14B across several compression ratios.
+Overall, CoSpaDi matches or improves upon the SVD-LLM baseline on most benchmarks and settings, indicating that sparse dictionary learning provides a stronger compression–quality trade-off than low-rank factorization beyond the standard evaluation suite.
+Table 6: Performance comparison of CoSpaDi vs sota SVD-LLM on Qwen3-8B and Qwen3-14B at different compression levels on different benchmarks. Best results are highlighted with bold.
+Model CR IFEval BBH MATH GPQA MUSR MMLU-Pro
+Qwen 3 8B 39.2 60.9 52.6 36.2 43.1 47.7
+SVDLLM 25.5 41.0 1.1 28.4 39.8 26.3
+CoSpaDi 0.2 28.9 45.3 2.0 28.6 42.1 31.5
+SVDLLM 22.9 34.4 1.0 25.6 41.4 18.8
+CoSpaDi 0.3 25.2 38.2 1.0 24.8 38.4 22.8
+SVDLLM 22.7 30.2 0.8 23.1 37.7 11.6
+CoSpaDi 0.4 26.1 32.7 0.8 26.0 38.1 16.7
+Qwen 3 14B 43.3 63.2 53.3 38.6 40.7 53.3
+SVDLLM 25.5 50.1 1.6 29.2 41.5 32.7
+CoSpaDi 0.2 29.6 54.2 4.2 30.6 40.7 38.9
+SVDLLM 23.0 34.3 1.1 26.4 39.7 12.8
+CoSpaDi 0.4 26.6 38.3 1.2 25.5 38.0 15.7
+A.7 K-SVD and Power Iteration Analysis
 We conducted ablation studies to assess the effect of the number of K-SVD iterations and power iterations on performance using Llama3.2-1B with fixed $\rho=2$.
@@
 Figure 8: Average benchmark accuracy and WikiText perplexity with respect to the number of K-SVD iterations (left) and the number of power iterations (right) for Llama3.2-1B with $\rho=2$
-A.9 Use of Large Language Models
-The authors acknowledge the use of a large language model (LLM) solely for language editing and grammatical refinement of the current manuscript.
-All scientific content, analysis, and interpretations presented herein are the sole responsibility of the authors.